ON SOME EXTENSIONS OF QUATERNIONS

By

William Rowan Hamilton

(Philosophical Magazine (4th series): vol. vii (1854), pp. 492–499, vol. viii (1854), pp. 125–137, 261–9, vol. ix (1855), pp. 46–51, 280–290.)

Edited by David R. Wilkins

NOTE ON THE TEXT

The paper On some Extensions of Quaternions, by Sir William Rowan Hamilton, appeared in 5 instalments in volumes vii–ix of The London, Edinburgh and Dublin Philosophical Magazine and Journal of Science (4th Series), for the years 1854–1855. Each instalment (including the last) ended with the words ‘To be continued’.

The articles of this paper appeared as follows:—

section i.  articles 1–6  Supplementary 1854 vol. vii (1854), pp. 492–499,
section ii. articles 7–16 August 1854  vol. viii (1854), pp. 125–137,
section iii. articles 17–25 October 1854  vol. viii (1854), pp. 261–269,
section iv. articles 26–29 January 1855  vol. ix (1855), pp. 46–51,
section v. articles 30–36 April 1855  vol. ix (1855), pp. 280–290.

(Articles 1–6 appeared in the supplementary number of the Philosophical Magazine which appeared in the middle of 1854.)

Some errata noted by Hamilton have been corrected.

The headings ‘Section I.’, ‘Section II.’ and ‘Section III.’ were not included for the first three sections of the original text, but analogous headings were included for the final two sections.

The paper On some Extensions of Quaternions, is included in The Mathematical Papers of Sir William Rowan Hamilton, vol. iii (Algebra), edited for the Royal Irish Academy by H. Halberstam and R. E. Ingram (Cambridge University Press, Cambridge, 1967).

David R. Wilkins Dublin, March 2000

Edition corrected—February 2001.

On some Extensions of Quaternions*. By Sir William Rowan Hamilton, LL.D., M.R.I.A., F.R.A.S., Corresponding Member of the French Institute, Hon. or Corr. Member of several other Scientific Societies in British and Foreign Countries, Andrews’ Professor of Astronomy in the University of Dublin, and Royal Astronomer of Ireland†.

[The London, Edinburgh and Dublin Philosophical Magazine and Journal of Science, 4th series, vol. vii (1854), pp. 492–99, vol. viii (1854), pp. 125–37, 261–9, vol. ix (1855), pp. 46–51, 280–90.]

Section I.

[1.] Conceive that in the polynomial expressions,

\begin{matrix}P & = ι_{0}x_{0} + ι_{1}x_{1} + … + ι_{n}x_{n} = Σιx, \\ P′ & = ι_{0}x′_{0} + ι_{1}x′_{1} + … + ι_{n}x′_{n} = Σιx′, \\ P′′ & = ι_{0}x′′_{0} + ι_{1}x′′_{1} + … + ι_{n}x′′_{n} = Σιx′′, \end{matrix} \}  (1)

the symbols x_{0} … x_{n}, which we shall call the constituents of the polynome P, and in like manner that the constituents x′_{0} … x′_{n} of P′, and x′′_{0} … x′′_{n} of P′′, are subject to all the usual rules of algebra, and to no others; but that the other symbols ι_{0} … ι_{n}, by which those constituents of each polynome are here symbolically multiplied, are not all subject to all those usual rules: and that, on the contrary, these latter symbols are subject, as a system, to some peculiar laws, of comparison and combination. More especially, let us conceive, in the first place, that these n + 1 symbols, of the form ι_{f}, are and must remain unconnected with each other by any linear relation, with ordinary algebraical coefficients; whence it will follow that an equality between any two polynomial expressions of the present class requires that all their corresponding constituents should be separately equal, or that

if P′ = P, then x′_{0} = x_{0}, x′_{1} = x_{1}, … x′_{n} = x_{n} :  (2)

and therefore, in particular, that the evanescence of any one such polynome P requires the vanishing of each constituent separately; so that

if P = 0, then x_{0} = 0, x_{1} = 0, … x_{n} = 0.  (3)

* See the work entitled, “Lectures on Quaternions,” by the present writer. (Hodges and Smith, Dublin, 1853.)

† Communicated by the Author.

In the second place, we shall suppose that all the usual rules of addition and subtraction extend to these new polynomes, and to their terms; and that the symbols ι, like the symbols x, are distributive in their operation; whence it will follow that

P′ ± P = ι_{0}(x′_{0} ± x_{0}) + … + ι_{n}(x′_{n} ± x_{n}),  (4)

or that

Σιx′ ± Σιx = Σι(x′ ± x) :  (5)

and as a further connexion with common algebra, we shall conceive that each separate symbol of the form ι may combine commutatively as a factor with each of the form x, and with every other algebraic quantity, so that ιx = xι, and that therefore the polynome P may be thus written,

P = x_{0}ι_{0} + x_{1}ι_{1} + … + x_{n}ι_{n} = Σxι.  (6)

But, third, instead of supposing that the symbols ι combine thus in general commutatively, among themselves, as factors or as operators, we shall distinguish generally between the two inverted (or opposite) products, ιι′ and ι′ι, or ι_{f}ι_{g} and ι_{g}ι_{f}; and shall conceive that all the (n + 1)^{2} binary products (ιι′), including squares (ι^{2} = ιι), of the n + 1 symbols ι, are defined as being each equal to a certain given or originally assumed polynome, of the general form (1), by (n + 1)^{2} equations of the following type,

ι_{f}ι_{g} = (fg0)ι_{0} + (fg1)ι_{1} + … + (fgh)ι_{h} + … + (fgn)ι_{n};  (7)

the (n+1)^{3} coefficients, or constituents, of the form (fgh), which we shall call the “constants of multiplication,” being so many given, or assumed, algebraic constants, of which some may vanish, and which we do not here suppose to satisfy generally the relation, (fgh) = (gfh). And thus the product of any two given polynomes, P and P′, of the form (1), combined in a given order as factors, becomes equal to a third given polynome, P′′, of the same general form,

P′′ = PP′ = Σx_{f}ι_{f} . Σx′_{g}ι_{g} = Σx′′_{h}ι_{h};  (8)

the summations extending still from 0 to n, and the constituent x′′_{h} of the product admitting of being thus expressed:

x′′_{h} = Σ(fgh)x_{f}x′_{g}.  (9)

As regards the subjection of the symbols ι to the associative law of multiplication, expressed by the formula,

ι . ι′ι′′ = ιι′ . ι′′,

we shall make no supposition at present.

[2.] As a first simplification of the foregoing very general* conception, let it be now supposed that

ι_{0} = 1;  (10)

* Some account of a connected conception respecting Sets, considered as including Quaternions, may be found in the Preface to the Lectures already cited. the n other symbols, ι_{1}, ι_{2}, … ι_{n}, being thus the only ones which are not subject to all the ordinary rules of algebra. Then because

ι_{0}ι_{g} = ι_{g}, ι_{f}ι_{0} = ι_{f},  (11)

it will follow that if either of the two indices f and g be = 0, the constant of multiplication (fgh) is either = 1 or = 0, according as h is equal or unequal to the other of those two indices; and we may write,

(0fh) = (f0h) = 0, if h > < f;  (12)
(0ff) = (f0f) = 1.  (13)

With this simplification, the number of the arbitrary or disposable constants of the form (fgh), which are not thus known already to have the value 0 or 1, is reduced from (n + 1)^{3} to (n + 1)n^{2}; because we may now suppose that f and g are each > 0, or that they vary only from 1 to n. For we may write,

P = p + ϖ, P′ = p′ + ϖ′,  (14)

where

\begin{matrix}p & = ι_{0}x_{0} = x_{0}, & ϖ & = ι_{1}x_{1} + … + ι_{f}x_{f} + … + ι_{n}x_{n}, \\ p′ & = ι_{0}x′_{0} = x′_{0}, & ϖ′ & = ι_{1}x′_{1} + … + ι_{g}x′_{g} + … + ι_{n}x′_{n}; \end{matrix} \}  (15)

and then, by observing that p and p′ are symbols of the usual and algebraical kind, shall have this expression for the product of two polynomes:

P′′ = PP′ = (p + ϖ)(p′ + ϖ′) = pp′ + pϖ′ + p′ϖ + ϖϖ′;  (16)

where the last term, or partial product, ϖϖ′, is now the only one for which any peculiar rules are required.

[3.] When the polynome P has thus been decomposed into two parts, p and ϖ, of which the one (p) is subject to all the usual rules of algebraical calculation, but the other (ϖ) to peculiar rules; and when these two parts are thus in such a sense heterogeneous, that an equation between two such polynomes resolves itself immediately into two separate equations, one between parts of the one kind, and the other between parts of the other kind; so that

if P = P′, or p + ϖ = p′ + ϖ′, then p = p, and ϖ = ϖ′;  (17)

we shall call the former part (p) the scalar part, or simply the scalar, of the polynome P, and shall denote it, as such, by the symbol S . P, or SP; and we shall call the latter part (ϖ) the vector part, or simply the vector, of the same polynome, and shall denote this other part by the symbol V.P, or VP: these names (scalar and vector), and these characteristics (S and V), being here adopted as an extension of the phraseology and notation of the Calculus of Quaternions*, in which such scalars and vectors receive useful geometrical interpretations.

* See Lectures, passim.

From the same calculus we shall here borrow also the conception and the sign of conjugation; and shall say that any two polynomes (such as those represented by p + ϖ and p − ϖ) are conjugate, if they have equal scalars (p), but opposite vectors (±ϖ): and if either of these two polynomes be denoted by P, then the symbol K.P, or KP, shall be employed to represent the other; K being thus used (as in quaternions) as the characteristic of conjunction. With these notations, and with the recent significations of p and ϖ,

p = S(p + ϖ), ϖ = V(p + ϖ), p − ϖ = K(p + ϖ);  (18)

or, writing P and P′ for p + ϖ and p − ϖ,

P′ = KP, if SP′ = SP, and VP′ = −VP;  (19)

and generally, for any polynome P, of the kind here considered,

P = SP + VP, KP = SP − VP.  (20)

We may also propose to call the n symbols ι_{1} … ι_{n} by the general name of vector-units, as the symbol ι_{0} has been equated in (10) to the scalar-unit, or to 1; and may call that equation (10) the unit-law, or more fully, the law of the primary unit.

[4.] Already, from these few definitions and notations, a variety of symbolical consequences can be deduced, which have indeed already occurred in the Calculus of Quaternions, but which are here taken with enlarged significations, and without reference to interpretation in geometry. For example, in the general equations (20), we may abstract from the operand, that is, from the polynome P, and may write more briefly (as in quaternions),

1 = S + V, K = S − V;  (21)

whence

S = \frac{1}{2}(1 + K), V = \frac{1}{2}(1 − K);  (22)

or more fully,

SP = \frac{1}{2}(P + P′), VP = \frac{1}{2}(P − P′), if P′ = KP.  (23)

Again, since (with the recent meanings of p and ϖ),

\begin{matrix}Sp = p, Vp = 0, Kp = p, Sϖ = 0, Vϖ = ϖ, Kϖ = −ϖ, \\ S(p − ϖ) = p, V(p − ϖ) = −ϖ, K(p − ϖ) = p + ϖ, \end{matrix} \}  (24)

we may write

\begin{matrix}SSP = SP, VSP = 0 = SVP, VVP = VP, \\ SKP = SP = KSP, VKP = −VP = KVP, KKP = P; \end{matrix} \}  (25)

or more concisely,

\begin{matrix}S^{2} = S, VS = SV = 0, V^{2} = V, \\ SK = KS = S, VK = KV = −V, K^{2} = 1. \end{matrix} \}  (26)

The operations, S, V, K are evidently distributive,

SΣ = ΣS, VΣ = ΣV, KΣ = ΣK;  (27)

and hence it is permitted to multiply together any two of the equations (21), (22), or to square any one of them, as if S, V, K were ordinary algebraical symbols, and the results must be found to be consistent with those equations themselves, and with the relations (26). Thus, squaring and multiplying the equations (21), we obtain

\begin{matrix}1^{2} & = (S + V)^{2} = S^{2} + V^{2} + 2SV = S + V = 1, \\ K^{2} & = (S − V)^{2} = S^{2} + V^{2} − 2SV = S + V = 1, \\ 1K & = (S + V)(S − V) = S^{2} − V^{2} = S − V = K; \end{matrix} \}  (28)

and the equations (22) give similarly,

\begin{matrix}S^{2} & = \frac{1}{4}(1 + K)^{2} = \frac{1}{4}(1 + K^{2} + 2K) = \frac{1}{2}(1 + K) = S; \\ V^{2} & = \frac{1}{4}(1 − K)^{2} = \frac{1}{4}(1 + K^{2} − 2K) = \frac{1}{2}(1 − K) = V; \\ SV & = VS = \frac{1}{4}(1 + K)(1 − K) = \frac{1}{4}(1 − K^{2}) = \frac{1}{4}(1 − 1) = 0. \end{matrix} \}  (29)

Again, if we multiply (22) by K, we get

\begin{matrix}KS & = \frac{1}{2}K(1 + K) = \frac{1}{2}(K + K^{2}) = \frac{1}{2}(K + 1) = S, \\ KV & = \frac{1}{2}K(1 − K) = \frac{1}{2}(K − K^{2}) = \frac{1}{2}(K − 1) = −V; \end{matrix} \}  (30)

all which results are seen to be symbolically true, and other verifications of this sort may easily be derived, among which the following may be not unworthy of notice:

(S ± V)^{2m} = 1, (S ± V)^{2m+1} = S ± V, (\frac{1 ± K}{2})^{m} = \frac{1 ± K}{2},  (31)

where m is any positive whole number.

[5.] As a second simplification of the general conception of polynomes of the form (1), which will tend to render the laws of their operations on each other still more analogous to those of the quaternions, let it be now conceived that the choice of the “constants of multiplication,” (fgh), is restricted by the following condition, which may be called the “Law of Conjugation:”

K . ιι′ = ι′ι, or K . ι_{f}ι_{g} = ι_{g}ι_{f};  (32)

namely the condition that, “opposite (or inverted) products of any two of the n symbols ι_{1}, … ι_{n}, shall always be conjugate polynomes.” The indices f and g being still supposed to be each > 0, the constants of multiplication (fgh), which had remained arbitrary and disposable in [2.], after that first simplification which consisted in supposing ι_{0} = 1, come now to be still further reduced in number, from (n + 1)n^{2} to \frac{1}{2}n(n^{2} + 1). For we have now, by operating with S on the equation (32), the following formula of relation between those constants,

(fg0) = (gf0);  (33)

and by comparing coefficients of ι_{h}, this other formula is obtained,

−(fgh) = (gfh), if h > 0;  (34)

whence

(ffh) = 0, if h > 0.  (35)

Writing, for conciseness,

(fg0) = (fg), (ff) = (f),  (36)

the squares, ι^{2}, of the n vector-units ι, will thus reduce themselves to so many constant scalars,

ι_{1}^{2} = (1), ι_{2}^{2} = (2), … ι_{f}^{2} = (f), … ι_{n}^{2} = (n);  (37)

and besides these, we shall have (n + 1) × \frac{n(n − 1)}{2} = \frac{1}{2}(n^{3} − n) other scalars, as constants of multiplication; namely the constituents (fgh) of the polynomial expansions of all the binary products ιι′ or ι_{f}ι_{g}, of unequal vector-units, taken in any one selected order, for instance so that g > f; it being unnecessary now, on account of the formulæ of relation (33), (34), to attend also to the opposite order of the two factors, if the object be merely to determine the number of the independent constants, which number is thus found to be n + \frac{1}{2}(n^{3} − n) = \frac{1}{2}(n^{3} + n), as above stated. Such then is the number of the constants of multiplication, including n of the form (f), and \frac{1}{2}n(n − 1) of the form (fg), besides others of the form (fgh), which remain still arbitrary, or disposable, after satisfying, first, the Unit-Law, ι_{0} = 1, and, second, the Law of Conjugation, K . ιι′ = ι′ι.

[6.] From this law of conjugation, (32), several general consequences follow. For, first, we see from it that “the square of every vector is a scalar,” which may be thus expanded:

\begin{matrix}ϖ & = (ι_{1}x_{1} + … + ι_{n}x_{n})^{2} = (1)x_{1}^{2} + (2)x_{2}^{2} + … + (n)x_{n}^{2} \\ & + 2(12)x_{1}x_{2} + 2(13)x_{1}x_{3} + … + 2(fg)x_{f}x_{g} + …; \end{matrix} \}  (38)

that is, more briefly,

(Σιx)^{2} = Σ(e)x_{e}^{2} + 2Σ(fg)x_{f}x_{g},  (39)

the summations extending to values of the indices > 0 and g being > f. In the second place, and more generally, “inverted products of any two vectors are equal to conjugate polynomes;” or in symbols,

ϖ′ϖ = K . ϖϖ′,  (40)

whatever two vectors may be denoted by ϖ and ϖ′. In fact, these two products have (according to the definition [3.] of conjugates) one common scalar part, but opposite vector parts,

\begin{matrix}S . ϖ′ϖ & = S . ϖϖ′ = Σ(e)x_{e}x′_{e} + Σ(fg)(x_{f}x′_{g} + x_{g}x′_{f}); \\ −V . ϖ′ϖ & = V . ϖϖ′ = Σ(fgh)(x_{f}x′_{g} − x_{g}x′_{f})ι_{h} : \end{matrix} \}  (41)

whence also we may write, as in quaternions,

S . ϖϖ′ = \frac{1}{2}(ϖϖ′ + ϖ′ϖ), V . ϖϖ′ = \frac{1}{2}(ϖϖ′ − ϖ′ϖ).  (42)

And, thirdly, the result (40) may be still further generalized as follows: “The conjugate of the product of any two polynomes is equal to the product of their conjugates, taken in an inverted order;” or in symbols,

K . PP′ = KP′ . KP.  (43)

In fact, we have now, by (16), (24), (27) and (40),

KP′′ = K . PP′ = K . (p + ϖ)(p′ + ϖ′)
= K(pp′ + pϖ′ + p′ϖ + ϖϖ′)
= pp′ − pϖ′ − p′ϖ + ϖ′ϖ
= (p′ − ϖ′)(p − ϖ) = KP′ . KP,  (44)

as asserted in (43). It follows also, fourthly, that “the product of any two conjugate polynomes is a scalar, independent of their order, and equal to the difference of the squares of the scalar and vector parts of either of them;” for,

if P′ = KP then PP′ = (p + ϖ)(p − ϖ) = p^{2} − ϖ^{2};  (45)

where ϖ^{2} is, by (38) or (39), a scalar. And if we agree to call the square root (taken with a suitable sign) of this scalar product of two conjugate polynomes, P and KP, the common tensor of each, and to denote it by the symbol TP; if also we give the name of versor to the quotient of a polynome divded by its own tensor, and denote this quotient by the symbol UP: we shall then be able to establish several general formulæ, as extensions from the theory of quaternions. For we shall have

TP = TKP = √(PKP) = \{(SP)^{2} − (VP)^{2}\}^{\frac{1}{2}} ;  (46)
T(p ± ϖ) = (p^{2} − ϖ^{2})^{\frac{1}{2}} ; Tp = (p^{2})^{\frac{1}{2}} , Tϖ = (−ϖ^{2})^{\frac{1}{2}} ;  (47)
UP = \frac{P}{√(PKP)}, U(p ± ϖ) = \frac{p ± ϖ}{(p^{2} − ϖ^{2})^{\frac{1}{2}}} ;  (48)
P = TP . UP = UP . TP;  (49)
TUP = UTP = 1; TTP = TP; UUP = UP :  (50)

with some other connected equations. But, although the chief terms (such as scalar, vector, conjugate, tensor, versor), and the main notations answering thereto (namely S, V, K, T, U), of the calculus of quaternions, along with several general formulæ resulting, come thus to receive extended significations, as applying to certain polynomial expressions which involve n vector-units, and for which as many as \frac{1}{2}(n^{3} + n) constants of multiplication are still left arbitrary and disposable; yet it must be observed, that we have not hitherto established any modular property of either of the two functions, which have been called above the tensor and versor of a polynome; nor any associative law, for the multiplication of three such polynomes together.

Observatory of Trinity College, Dublin, June 6, 1854.

Section II.

[7.] Let us now consider generally the associative law of multiplication, which may be expressed by the formula already mentioned but reserved in [1.],

ι . ι′ι′′ = ιι′ . ι′′;  (51)

or by this other equation,

ι_{e} . ι_{f}ι_{g} = ι_{e}ι_{f} . ι_{g} :  (52)

and let us inquire into the conditions under which this law shall be fulfilled, for any 3 unequal or equal symbols of the form ι.

If the conception of the polynomial expression

P = Σιx = ι_{0}x_{0} + ι_{1}x_{1} + … + ι_{n}x_{n}, (1)

be no further restricted than it was in [1.], then each of the three indices e, f, g, in the equation (52), may receive any one of the n + 1 values from 0 to n; so that there are in this case (n + 1)^{3} associative conditions of this form (52), whereof each, by comparison of the coefficients of the n + 1 symbols ι, breaks itself up into n + 1 separate equations, of the ordinary algebraical kind, making in all no fewer than (n + 1)^{4} algebraical relations, to be satisfied, if possible, by the (n+1)^{3} constants of multiplication, of the form (fgh): respecting which constants, it will be remembered that the general formula has been established,

ι_{f}ι_{g} = (fg0)ι_{0} + … + (fgh)ι_{h} + … + (fgn)ι_{n}. (7)

We may therefore substitute, in (52), the expressions,

\begin{matrix}ι_{f}ι_{g} & = Σ_{h}(fgh)ι_{h}, & ι_{e}ι_{f} & = Σ_{h}(efh)ι_{h}, \\ ι_{e}ι_{h} & = Σ_{k}(ehk)ι_{k}, & ι_{h}ι_{g} & = Σ_{k}(hgk)ι_{k}; \end{matrix} \}  (53)

and then, by comparing coefficients of ι_{k}, this associative formula (52) breaks itself up, as was just now remarked, into (n + 1)^{4} equations between the (n + 1)^{3} constants, which are all included in the following*:

Σ_{h}(fgh)(ehk) = Σ_{h}(efh)(hgk);  (54)

where the four indices e f g k may each separately receive any one of the n + 1 values from 0 to n, and the summations relatively to h are performed between the same limits.

* This formula (54) may be deduced from the equation (214) in p. 239 of the writer’s “Researches respecting Quaternions”, published in the Transactions of the Royal Irish Academy, vol. xxi, part 2, by changing there the letters r s t r′ s′ to f h g e k, and substituting the symbol (fgh) for n_{g,f,h}. Or the same formula (54) may be derived from one given in page (30) of the Preface to the same author’s Lectures on Quaternions, (Dublin, Hodges and Smith, 1853), by writing g f e k instead of f g g′ h′, and changing each of the two symbols 1_{g,f,h}, 1′_{g,f,h}, to (fgh). But the general reductions of the present paper have not been hitherto published.

[8.] Introducing next the simplification (10) of article [2.], or supposing ι_{0} = 1, which has been seen to reduce the number of the constants of multiplication from (n + 1)^{3} to (n + 1)n^{2}, we find that the number of equations to be satisfied by them is reduced in a still greater ratio, namely from (n + 1)^{4} to (n + 1)n^{3}. For, if we suppose the index g to become 0, and observe that each of the constants (f0h) and (0fh) is equal, by (12) and (13), to 0 or to 1, according as h is unequal or equal to f, we shall see that the sum in the left-hand member of the formula (54) reduces itself to the term (efk): but such is also in this case the value of the right-hand sum in the same formula, because in calculating that sum we need attend only to the value h = k, if g be still = 0. In like manner, if f = 0, each sum reduces itself to (egk); and if e = 0, the two sums become each = (fgk). If then any one of these three indices, e, f, g, be = 0, the formula (54) is satisfied: which might indeed have been foreseen, by observing that, in each of these three cases, one factor of each member of the equation (52) becomes = 1. We may therefore henceforth suppose that each of the three indices, e, f, g, varies only from 1 to n, or that

e > 0, f > 0, g > 0;  (55)

while k may still receive any value from 0 to n, and h still varies in the summations between these latter limits: and thus the number of equations, supplied by the formula (54), between the constants (fgh), is reduced, as was lately stated, to (n + 1)n^{3}; while the number of those constants themselves had been seen to be reduced to (n + 1)n^{2}, by the same supposition ι_{0} = 1.

[9.] Additional reductions are obtained by introducing the law of conjugation (32), or by supposing K . ι_{f}ι_{g} = ι_{g}ι_{f}, with the consequences already deduced from that law or equation in [5.]. Using Σ′ to denote a summation relatively to h from 1 to n, and taking separately the two cases where k = 0 and where k > 0, we have, for the first case, by (54),

Σ′(efh)(gh) = Σ′(fgh)(eh);  (56)

and for the second case,

(ef)(g0k) − (fg)(e0k) = Σ′{(efh)(ghk) + (fgh)(ehk)}.  (57)

No new conditions would be obtained by interchanging e and g; but if we cyclically change efg to fge, each of the two sums (56) is seen to be equal to another of the same form; and two new equations are obtained from (57), by adding which thereto we find,

0 = Σ′{(efh)(ghk) + (fgh)(ehk) + (geh)(fhk)};  (58)

and therefore,

(fg)(e0k) − (ef)(g0k) = Σ′(geh)(fhk).  (59)

When e = f, the equations (56) and (59) become, respectively,

0 = Σ′(fh)(fgh),  (60)

and

(fg)(f0k) − (f)(g0k) = Σ′(gfh)(fhk);  (61)

which are identically satisfied, if we suppose also f = g; the properties [5.] of the symbols (fgh) being throughout attended to: while, by the earlier properties [2.], the symbol (e0k) or (0ek) is equal to 0 or to 1, according as e and k are unequal or equal to each other. And no equations distinct from these are obtained by supposing e = g, or f = g, in (56) and (59). The associative conditions for which k = 0 are, therefore, in number, n(n − 1) of the form (60), and \frac{1}{3}n(n−1)(n−2) of the form (56); or \frac{1}{3}(n^{3} −n) in all. And the other associative conditions, for which k > 0, are, in number, n^{2}(n − 1) of the form (61), and \frac{1}{2}n^{2}(n − 1)(n − 2) of the form (59), or \frac{1}{2}(n^{4} − n^{3}) in all. It will, however, be found that this last number admits of being diminished by \frac{1}{2}(n^{2} −n), namely by one for each of the symbols of the form (fg); and that if, before or after this reduction, the associative equations for which k > 0 be satisfied, then those other \frac{1}{3}(n^{3} − n) conditions lately mentioned, for which k = 0, are satisfied also, as a necessary consequence. The total number of the equations of association, included in the formula (54), will thus come to be reduced to

\frac{1}{2}(n^{4} − n^{3}) − \frac{1}{2}(n^{2} − n), or to \frac{1}{2}n(n − 1)(n^{2} − 1);

but it may seem unlikely that even so large a number of conditions as this can be satisfied generally, by the \frac{1}{2}n(n^{2} + 1) constants of multiplication [5.]. Yet I have found, not only for the case n = 2, in which we have thus 5 constants and 3 equations, but also for the cases n = 3 and n = 4, for the former of which we have 15 constants and 24 equations, while for the latter we have 34 constants and 90 equations, that all these associative conditions can be satisfied: even in such a manner as to leave some degree of indetermination in the results, or some constants of multiplication disposable.

[10.] Without expressly introducing the symbols (fgh), results essentially equivalent to the foregoing may be deduced in the following way, with the help of the characteristics [3.] of operation, S, V, K. The formula of association (51) may first be written thus*:

ι Sι′ι′′ + ιVι′ι′′ = Sιι′ . ι′′ + Vιι′ . ι′′;  (62)

in which the symbols Sιι′ and Vιι′ are used to denote concisely, without a point interposed, the scalar and vector parts of the product ιι′, but a point is inserted, after those symbols, and before ι′′, in the second member, as a mark of multiplication: so that, in this abridged notation, Sιι′ . ι′′ and Vιι′ . ι′′ denote the products which might be more fully expressed as (S . ιι′) × ι′′ and (V . ιι′) × ι′′; while it has been thought unnecessary to write any point in the first member, where the factor ι occurs at the left hand. Operating on (62) by S and V, we find the two following equations of association, which are respectively of the scalar and vector kinds:

S(ι Vι′ι′′ − ι′′ Vιι′) = 0;  (63)

* There is here a slight departure from the notation of the Lectures on Quaternions, by the suppression of certain points, which circumstance in the present connexion cannot produce ambiguity.

V(ι Vι′ι′′ + ι′′ Vιι′) = ι′′ Sιι′ − ι Sι′ι′′;  (64)

because the law (32) of conjugation, ι′ι = Kιι′, gives, by (41),

Sϖ′ϖ = +Sϖϖ′, Vϖ′ϖ = −Vϖϖ′.

For the same reason, no essential change is made in either of the two equations, (63), (64), by interchanging ι and ι′′; but if we cyclically permute the three vector-units, ι ι′ ι′′, then (63) gives

S(ι Vι′ι′′) = S(ι′ Vι′′ι) = S(ι′′ Vιι′);  (65)

and there arise three equations of the form (64), which give, by addition,

V(ι Vι′ι′′ + ι′ Vι′′ι + ι′′ Vιι′) = 0;  (66)

and therefore conduct to three other equations, of the form*

V(ι Vι′ι′′) = ι′′ Sιι′ − ι′ Sι′′ι.  (67)

Equating ι′′ to ι, the two equations (65) reduce themselves to the single equation,

S(ι Vιι′) = 0;  (68)

and the formula (67) becomes

V(ι Vιι′) = ι^{2}ι′ − ιSιι′ :  (69)

both which results become identities, when we further equate ι′ to ι. And no equations of condition, distinct from these, are obtained by supposing ι′′ = ι′, or ι′ = ι, in (65) and (67). The number of the symbols ι being still supposed = n, and therefore by [5.] the number of the constants which enter into the expressions of their n^{2} binary products (including squares) being = \frac{1}{2}(n^{3} + n), these constants are thus (if possible) to be made to satisfy \frac{1}{3}(n^{3} − n) associative and scalar equations of condition, obtained through (63), from the comparison of the scalar parts of the two ternary products, ι . ι′ι′′ and ιι′ . ι′′; namely, n(n − 1) scalar equations of the form (68) and \frac{1}{3}n(n − 1)(n − 2) such equations, of the forms (65). And the same constants of multiplication must also (if the associative law is to be fulfilled) be so chosen as to satisfy \frac{1}{2}(n^{3} − n^{2}) vector equations, equivalent each to n scalar equations, or in all to \frac{1}{2}(n^{4} − n^{3}) scalar conditions, obtained through (64) from the comparison of the vector parts of the same two ternary products (51); namely, n(n − 1) vector equations of the form (69), and \frac{1}{2}n(n − 1)(n − 2) other vector equations, included in the formula (64). This new analysis therefore confirms completely the conclusion of the foregoing paragraph respecting the general existence of \frac{1}{2}(n^{4} − n^{3}) + \frac{1}{3}(n^{3} − n) associative and scalar equations of condition, between the \frac{1}{2}(n^{3} + n) disposable constants of multiplication, when the general conception of the polynomial expression P of [1.] is modified by the suppositions, ι_{0} = 1 in [2.], and ι′ι = Kιι′ in [5.]. At least the analysis of the present paragraph [10.] confirms what has been lately proved in [9.], that the number of conditions of association can be reduced so far; but the same analysis will also admit of being soon applied, so as to assist in proving the existence of those additional and general reductions which have been lately mentioned without proof, and which depress the number of conditions to be satisfied to \frac{1}{2}(n^{4} − n^{3}) − \frac{1}{2}(n^{2} − n). Meanwhile it may be useful to exemplify briefly the foregoing general reasonings for the cases n = 2, n = 3, that is, for trinomial and quadrinomial polynomes.

* This formula is one continually required in calculating with quaternions (compare page li of the Contents, prefixed to the author’s Lectures).

[11.] For the case n = 2, the two distinct symbols of the form ι may be denoted simply by ι and ι′; and the equations of association to be satisfied are all included in these two,

ι . ιι′ = ι^{2}ι′, ι′ . ι′ι = ι^{′2}ι;  (70)

which give, when we operate on them by S and V, two scalar equations of the form (68), and two vector equations of the form (69), equivalent on the whole to six scalar equations of condition, between the five constants of multiplication, (1) (2) (12) (121) (122), if we write, on the plan of the preceding articles,

ι^{2} = (1), ι^{′2} = (2), Sιι′ = (12), Vιι′ = −Vι′ι = (121)ι + (122)ι′.  (71)

From (68), or from (60), or in so easy a case by more direct and less general considerations, we find that the comparison of the scalar parts of the products (70) conducts to the two equations,

0 = (121)(1) + (122)(12) = (122)(2) + (121)(12).  (72)

From (69), or (61), we find that the comparison of the vector parts of the same products (70) gives immediately four scalar equations, which however are seen to reduce themselves to the three following:

(121)(122) = −(12); (122)^{2} = (1); (121)^{2} = (2);  (73)

the first of these occurring twice. And it is clear that the equations (72) are satisfied, as soon as we assign to (1), (2) and (12) the values given by (73). If then we write, for conciseness,

(121) = a, (122) = b,  (74)

we shall have, for the present case (n = 2), the values,

(1) = b^{2}, (2) = a^{2}, (12) = −ab.  (75)

And hence, (writing κ instead of ι′,) we see that the trinome*,

P = z + ιx + κy,  (76)

where x y z are ordinary variables, will possess all the properties of those polynomial expressions which have been hitherto considered in this paper, and especially the associative property, if we establish the formula of multiplication,

(ιx + κy)(ιx′ + κy′) = (bx − ay)(bx′ − ay′) + (aι + bκ)(xy′ − yx′);  (77)

wherein a and b are any two constants of the ordinary and algebraical kind. In this trinomial system,

z′′ + ιx′′ + κy′′ = (z + ιx + κy)(z′ + ιx′ + κy′),  (78)

* I am not aware that this trinomial expression (76), with the formula of multiplication (77), coincides with any of the triplet-forms of Professor De Morgan, or of Messrs. John and Charles Graves: but it is given here merely by way of illustration. if

\begin{matrix}x′′ & = zx′ + z′x + a(xy′ − yx′), \\ y′′ & = zy′ + z′y + b(xy′ − yx′), \\ z′′ & = zz′ + (bx − ay)(bx′ − ay′); \end{matrix} \}  (79)

we have therefore the two modular relations,

\begin{matrix}z′′ + bx′′ − ay′′ & = (z + bx − ay)(z′ + bx′ − ay′), \\ z′′ − bx′′ + ay′′ & = (z − bx + ay)(z′ − bx′ + ay′); \end{matrix} \}  (80)

that is to say, the functions z ± (bx − ay) are two linear moduli of the system. A general theory with which this result is connected will be mentioned a little further on. Geometrical interpretations (of no great interest) might easily be proposed, but they would not suit the plan of this communication.

[12.] For the case n = 3, or for the quadrinome

P = x_{0} + ι_{1}x_{1} + ι_{2}x_{2} + ι_{3}x_{3},  (81)

we may assume

\begin{matrix}ι_{1}^{2} & = a_{1}, & ι_{2}^{2} & = a_{2}, & ι_{3}^{2} & = a_{3}, \\ Sι_{2}ι_{3} & = b_{1}, & Sι_{3}ι_{1} & = b_{2}, & Sι_{1}ι_{2} & = b_{3}, \end{matrix} \}  (82)

and

\begin{matrix}Vι_{2}ι_{3} & = −Vι_{3}ι_{2} = ι_{1}l_{1} + ι_{2}m_{3} + ι_{3}n_{2}, \\ Vι_{3}ι_{1} & = −Vι_{1}ι_{3} = ι_{2}l_{2} + ι_{3}m_{1} + ι_{1}n_{3}, \\ Vι_{1}ι_{2} & = −Vι_{2}ι_{1} = ι_{3}l_{3} + ι_{1}m_{2} + ι_{2}n_{1}; \end{matrix} \}  (83)

and then the \frac{1}{2}(n^{4} − n^{3}) = 27 scalar equations of condition, included in the vector form,

V(ι . ι′ι′′) = V(ιι′ . ι′′),  (84)

are found on trial to reduce* themselves to 24; which, after elimination of the 6 constants of the forms here denoted by a and b, or previously by (f) and (fg), furnish 18 equations of condition between the 9 other constants, of the forms here marked l, m, n, or previously (fgh); and these 18 equations may be thus arranged†:

\begin{matrix}0 & = l_{1}(n_{1} − m_{1}) = l_{2}(n_{2} − m_{2}) = l_{3}(n_{3} − m_{3}), \\ 0 & = l_{2}(n_{1} − m_{1}) = l_{3}(n_{2} − m_{2}) = l_{1}(n_{3} − m_{3}), \\ 0 & = l_{3}(n_{1} − m_{1}) = l_{1}(n_{2} − m_{2}) = l_{2}(n_{3} − m_{3}); \end{matrix} \}  (85)

* The reason of this reduction is exhibited by the general analysis in [14.].

† For it is found that each of the three constants (eff) + (fgg) must give a null product, when it is multiplied by any one of the constants (e′f′g′), or by any one of these other constants (e′′f′′f′′) − (e′′g′′g′′); if each of the three systems, efg, e′f′g′, e′′f′′g′′, represent, in some order or other, but not necessarily in one common order, the system of the three unequal indices, 1, 2, 3.

\begin{matrix}0 & = n_{1}^{2} − m_{1}^{2} = n_{2}^{2} − m_{2}^{2} = n_{3}^{2} − m_{3}^{2}, \\ 0 & = (n_{2} + m_{2})(n_{1} − m_{1}) = (n_{3} + m_{3})(n_{2} − m_{2}) = (n_{1} + m_{1})(n_{3} − m_{3}), \\ 0 & = (n_{3} + m_{3})(n_{1} − m_{1}) = (n_{1} + m_{1})(n_{2} − m_{2}) = (n_{2} + m_{2})(n_{3} − m_{3}); \end{matrix} \}  (86)

they are therefore satisfied, without any restriction on l_{1} l_{2} l_{3}, by our supposing

n_{1} = m_{1}, n_{2} = m_{2}, n_{3} = m_{3};  (87)

but if we do not adopt this supposition, they require us to admit this other system of equations,

0 = l_{1} = l_{2} = l_{3} = n_{1} + m_{1} = n_{2} + m_{2} = n_{3} + m_{3}.  (88)

Whichever of these two suppositions, (87), (88), we adopt, there results a corresponding system of values of the six recently eliminated constants, of the forms a and b, or (f) and (fg); and it is found* that these values satisfy, without any new supposition being required, the \frac{1}{3}(n^{3} − n) = 8 scalar equations, included in the general form

S(ι . ι′ι′′) = S(ιι′ . ι′′),  (89)

which are required for the associative property.

[13.] In this manner I have been led to the two following systems of associative quadrinomials, which may be called systems (A) and (B); both possessing all those general properties of the polynomial expression P, which have been considered in the present paper; and one of them including the quaternions.

For the system (A), the quadrinomial being still of the form (81), or of the following equivalent form,

Q = w + ιx + κy + λz,  (90)

where w x y z are what were called in [1.] the constituents, the laws of the vector-units ι κ λ are all included in this formula of multiplication for any two vectors, such as

ρ = ιx + κy + λz, ρ′ = ιx′ + κy′ + λz′ :  (91)

(A) …  ρρ′ = (m_{1}^{2} − l_{2}l_{3})xx′ + (l_{1}m_{1} − m_{2}m_{3})(yz′ + zy′)
+ (m_{2}^{2} − l_{3}l_{1})yy′ + (l_{2}m_{2} − m_{3}m_{1})(zx′ + xz′)
+ (m_{3}^{2} − l_{1}l_{2})zz′ + (l_{3}m_{3} − m_{1}m_{2})(xy′ + yx′)
+ (ιl_{1} + κm_{3} + λm_{2})(yz′ − zy′)
+ (κl_{2} + λm_{1} + ιm_{3})(zx′ − xz′)
+ (λl_{3} + ιm_{2} + κm_{1})(xy′ − yx′);  (92)

* This fact of calculation is explained by the general analysis of [15.]. The values of a and b may be deduced from the formulæ, a_{1} = m_{1}^{2} −l_{2}l_{3}, b_{1} = l_{1}m_{1} −m_{2}n_{3}, with others cyclically formed from these. and it is clear that Quaternions* are simply that particular case of such quadrinomes (A), for which the six arbitrary constants l_{1} … m_{3} and the three vector-units ι κ λ receive the following values:

l_{1} = l_{2} = l_{3} = 1, m_{1} = m_{2} = m_{3} = 0, ι = i, κ = j, λ = k.  (93)

For the other associative quadrinomial system (B), which we may call for distinction tetrads, if we retain the expressions (90), (91), we must replace the formula of vectormultiplication (92) by one of the following form:

(B) …  ρρ′ = (lx + my + nz)(lx′ + my′ + nz′)
+ (κn − λm)(yz′ − zy′) + (λl − ιn)(zx′ − xz′)
+ (ιm − κl)(xy′ − yx′);  (94)

involving thus only three arbitrary constants, l m n, besides the three vector-units, ι κ λ; and apparently having no connexion with the quaternions, beyond the circumstance that one common analysis [12.] conducts to both the quadrinomes (A) and the tetrads (B).

As regards certain modular properties of these two quadrinomial systems, we shall shortly derive them as consequences of the general theory of polynomes of the form P, founded on the principles of the foregoing articles.

[14.] In general, the formula (59) gives, by [2.], the two following equations, which may in their turn replace it, and are, like it, derived from the comparison of the vector parts of the general associative formula, or from the supposition that k > 0 in (54):

(fg) = Σ(geh)(fhe), if e > < g;  (95)
0 = Σ(geh)(fhk), if k > < e, k > < g;  (96)

the summation extending in each from h = 1 to h = n. Interchanging f and g in (95), we have

(gf) = Σ(feh)(ghe), if e > < f;  (97)

and making g = f, in either (95) or (97), we obtain the equation,

(f) = Σ(feh)(fhe), if e > < f.  (98)

For each of the n symbols (f), there are n − 1 distinct expressions of this last form, obtained by assigning different values to e; and when these expressions are equated to each other, there result n(n − 2) equations between the symbols of the form (fgh). For each of the \frac{1}{2}n(n − 1) symbols of the form (fg), where f and g are unequal, there are n − 1 expressions (95), and n − 1 other expressions of the form (97), because, by (33) and (36), (gf) = (fg); and thus it might seem that there should arise, by equating these 2n−2 expressions for each symbol (fg),

* See the author’s Lectures, or the Philosophical Magazine for July, 1844, in which the first printed account of quaternions was given. as many as 2n − 3 equations from each, or \frac{1}{2}n(n − 1)(2n − 3) equations in all between the symbols (fgh). But if we observe that the sums of the n − 1 expressions (95) for (fg), and of the n − 1 expressions (97) for (gf) are, respectively,

(n − 1)(fg) = Σ_{e}Σ_{h}(geh)(fhe), (n − 1)(gf) = Σ_{e}Σ_{h}(feh)(ghe);  (99)

where the summations may all be extended from 1 to n, because (ffh) and (ggh) are each = 0, by (35), since h > 0; and that these two double sums (99) are equal; we shall see that the formula

(gf) = (fg),  (100)

though true, gives no information respecting the symbols (fgh): or is not to be counted as a new and distinct equation, in combination with the n − 1 equations (95), and the n − 1 equations (97). In other words, the comparison of the sums (99) shows that we may confine ourselves to equating separately to each other, for each pair of unequal indices f and g, the n − 1 expressions (95) for the symbol (fg), and the n − 1 other expressions (97) for the symbol (gf), without proceeding afterwards to equate an expression of the one set to an expression of the other set. We may therefore suppress, as unnecessary, an equation of the form (100), for each of the \frac{1}{2}n(n − 1) symbols of the form (fg), or for each pair of unequal indices f and g, as was stated by anticipation towards the close of paragraph [9.]. There remain, however, 2(n − 2) equations of condition, between the symbols (fgh), derived from each of those \frac{1}{2}n(n − 1) pairs; or as many as n(n − 1)(n − 2) equations in all, obtained in this manner from (95) and (97), regarded as separate formulæ. Thus, without yet having used the formula (96), we obtain, with the help of (98), by elimination of the symbols (f), (fg), (gf), through the comparison of n − 1 expressions for each of those n^{2} symbols, n^{2}(n − 2) equations of condition, homogeneous and of the second dimension, between the symbols of the form (fgh). And without any such elimination, the formula (96) gives immediately \frac{1}{2}n^{2}(n − 1)(n − 2) other equations of the same kind between the same set of symbols; because after choosing any pair of unequal indices e and g, we may combine this pair with any one of the n values of the index f, and with any one of the n−2 values of k, which are unequal both to e and to g. There are therefore, altogether, \frac{1}{2}n^{2}(n + 1)(n − 2) homogeneous equations of the second dimension, obtained by comparison of the vector parts of the general formula of association, to be satisfied by the \frac{1}{2}n^{2}(n − 1) symbols of the form (fgh).

[15.] To prove now, generally, that when the vector parts of the associative formula are thus equal, the scalar parts of the same formula are necesarily equal also, or that the system of conditions (56) in [9.] is included in the system (57) or (59); we may conveniently employ the notations S and V, and pursue the analysis of paragraph [10.], so as to show that the system of equations (65), including (68), results from the system (67), including (69); or that if the formula (84) be satisfied for every set of three unequal or equal vector-units, ι ι′ ι′′, then, for every such set, the formula (89) is satisfied also. For this purpose, I remark that the formula of vector-association (67), when combined with the distributive principle of multiplication [1.], and of operation with S and V [5.], gives generally, as in quaternions, the transformation

Vρ Vστ = τ Sρσ − σ Sρτ;  (101)

where ρ, σ, τ may denote any three vectors, and the symbol Vρ Vστ is used to signify concisely the vector part of the product ρ×V(στ); whence also we may derive by (41) this other general transformation,

V(Vστ . ρ) = σ Sρτ − τ Sρσ.  (102)

If then we write

Vστ = ρ′, Vτρ = σ′, Vρσ = τ′,  (103)

and introduce another arbitrary vector ϖ, we shall have

Vρ′ϖ = σ Sτϖ − τ Sσϖ;  (104)

and therefore

Vρ Vρ′ϖ = τ′ Sτϖ + σ′ Sσϖ;  (105)

but also

Vρ Vρ′ϖ = ϖ Sρρ′ − ρ′ Sρϖ;  (106)

whence

ϖ Sρρ′ = ρ′ Sρϖ + σ′ Sσϖ + τ′ Sτϖ,  (107)

and consequently

Sρρ′ = Sσσ′ = Sττ′;  (108)

but this is precisely by (103) the formula of scalar-association (65), stated in its most general form. The general dependence of (65) on (67), or of (56) on (57), is therefore proved to exist; and the \frac{1}{3}(n^{3} − n) associative conditions, for which k = 0 in (54), are seen to be consequences of \frac{1}{2}(n^{4} − n^{3}) other conditions for which k > 0; or even of those conditions diminished in number by \frac{1}{2}(n^{2} − n), according to what was stated by anticipation in [9.], and has been proved by the analysis of [14.]. This result is the more satisfactory, because otherwise the conditions of association would essentially involve a system of homogeneous equations of the third dimension relatively to the symbols (fgh), obtained by substituting in (56) the expressions (95) or (97) for the symbols of the form (fg), including the values (98) of the symbols (f). But we see now (as above stated) that the total number of distinct conditions may be reduced to \frac{1}{2}(n^{4} − n^{3}) − \frac{1}{2}(n^{2} − n), between the total number \frac{1}{2}(n^{3} + n) of constants of multiplication; or finally, after the elimination of the \frac{1}{2}(n^{2} + n) symbols of the forms (f) and (fg), to a system of homogeneous equations of the second dimension, namely, those determined in [14.], of which the number amounts (as in that paragraph) to

\frac{1}{2}(n^{4} − n^{3}) − n^{2} = \frac{1}{2}n^{2}(n + 1)(n − 2),  (109)

between the symbols (fgh), whereof the number is

\frac{1}{2}(n^{3} + n) − \frac{1}{2}(n^{2} + n) = \frac{1}{2}n^{2}(n − 1).  (110)

[16.] For example, when n = 2, the two constants (121) and (122) have been seen in [11.] to be unresticted by any condition. When n = 3, we have 9 constants, lately denoted by l_{1} l_{2} l_{3} m_{1} m_{2} m_{3} n_{1} n_{2} n_{3}, wherewith to satisfy 18 homogeneous equations of the second dimension, namely those marked (85) and (86) in [12.]; which it has been seen to be possible to do, in two distinct ways (A) and (B), and even so as to leave some of the constants arbitrary, in each of the two resulting systems, of associative quadrinomes and tetrads. A similar result has been found by me to hold good for the case n = 4, or for the case of associative quines, such as

P = w + ιx + κy + λz + μu,  (111)

involving four vector-units ι κ λ μ, which obey the laws of conjugation (32), and of association (51). For although there are in this case only 24 = \frac{1}{2}n^{2}(n − 1) constants of the form (fgh), to satisfy 80 = \frac{1}{2}n^{2}(n + 1)(n − 2) homogeneous equations of the second dimension, yet I have found that the forms* of these equations are such as to allow this to be done in various ways, and even without entirely determining the constants. And it appears not impossible that similar results may be obtained for higher values of n; or that associative† polynomes of higher orders than quines may be discovered.

Observatory of Trinity College, Dublin, July 4, 1854.

* The subject may be illustrated by the very simple remark, that although the four equations tx = 0, ty = 0, ux = 0, uy = 0, are such that no three of them include the fourth, since we might (for example) satisfy the three first alone by supposing t = 0, x = 0, yet they can all four be satisfied together by supposing either x = 0, y = 0, or t = 0, u = 0. Compare the equations (85) or (86), which are of the forms tx = 0, ty = 0, tz = 0, ux = 0, uy = 0, uz = 0, vx = 0, vy = 0, vz = 0. In the theory of quines, however, the forms are not quite so simple.

† The octaves, or octonomial expressions, which Mr. Cayley published in the Philosophical Magazine for March 1845, and which had been previously but privately communicated to me by Mr. J. T. Graves about the end of 1843, after my communication to him of the quaternions, are not associative polynomes. Thus in Mr. Cayley’s notation, the four following of his seven types, (123) (624) (176) (734), give ι_{1} . ι_{2}ι_{4} = ι_{1}ι_{6} = −ι_{7}, but ι_{1}ι_{2} . ι_{4} = ι_{3}ι_{4} = +ι_{7}; or with Mr. Graves’s symbols, the triads i j k, i o n, j l n, k l o, give i.jl = in = −o, but ij.l = kl = +o. See note to page (61) of the Preface to my Lectures. It was my perceiving this latter property of Mr. Graves’s symbols in 1844, which chiefly discouraged me from pursuing the study of those octaves, as a species of extensions of the quaternions, which Mr. Graves as well as Mr. Cayley had designed them to be, and which in one sense no doubt they are.

Section III.

[17.] The following remarks may be useful, as serving to illustrate and develope the general analysis contained in some of the preceding paragraphs, especially in [14.] and as adapted to give some assistance towards any future study of associative polynomes, such as quines, of an order higher than quadrinomes, but subject like them to the law of conjugation (32).

The expression (98) may be thus more fully written:

(f) = (fee)^{2} + (feg)(fge) + Σ^{‵}(feh)(fhe);  (112)

where e f g h are all supposed to be unequal; the summation Σ^{‵} being performed relatively to h, for all those n − 3 values of the latter, which are distinct from each of the three former indices. Interchanging e and g, and subtracting, we eliminate the symbol (f), and obtain the following formula:

I. …  (fee)^{2} − (fgg)^{2} = Σ^{‵}{(fgh)(fhg) − (feh)(fhe)};  (113)

which type I. includes generally n(n − 2) distinct and homogeneous equations, of the second dimension, with 2(n − 2) terms in each, between the \frac{1}{2}n^{2}(n − 1) symbols of the form (fgh). Thus, for the case of quadrinomials (n = 3), by writing, in agreement with (82) and (83),

a_{1} = (1), b_{1} = (23), l_{1} = (231), m_{1} = (313), n_{1} = (122),  (114)

and suppressing the sum Σ^{‵}, we have by (112) the two expressions (compare a note to [12.]):

a_{1} = m_{1}^{2} − l_{2}l_{3} = n_{1}^{2} − l_{2}l_{3};  (115)

together with four others formed from these, by cyclical permutation of the indices 1, 2, 3; and we are thus conducted, by elimination of the three symbols a_{1}, a_{2}, a_{3}, to three equations of the form n_{1}^{2} = m_{1}^{2}; that is, to the 3 equations on the first line of (86), involving each 2 terms. For quines (n = 4), if we make also, with the same permitted permutations,

\begin{matrix}a_{4} = (4), c_{1} = (14), p_{1} = (234), \\ r_{1} = (141), s_{1} = (142), t_{1} = (143), u_{1} = (144), \end{matrix} \}  (116)

the index h receives one value under each sign of summation Σ^{‵}, and the resulting formulæ may be thus written:

(a_{1} + l_{3}l_{2} + p_{2}t_{1} − s_{1}p_{3} = ) n_{1}^{2} + p_{2}t_{1} = m_{1}^{2} − s_{1}p_{3} = u_{1}^{2} + l_{3}l_{2};  (117)
(a_{4} − s_{1}t_{2} − s_{2}t_{3} − s_{3}t_{1} = ) r_{1}^{2} − s_{2}t_{3} = r_{2}^{2} − s_{3}t_{1} = r_{3}^{2} − s_{1}t_{2};  (118)

where the line (117) is equivalent to three lines of the same form: so that the elimination of a_{1} … a_{4} conducts here to 8 equations, of 4 terms each, between the 24 symbols of the form (fgh), or l_{1} … u_{3}, as by the general theory it ought to do. For polynomes of higher orders (n > 4), we have the analogous equations,

(f) − (feg)(fge) − (fgk)(fkg) − (fke)(fek)
= (fee)^{2} − (fgk)(fkg) + Σ^{‵}(feh)(fhe)
= (fgg)^{2} − (fke)(fek) + Σ^{‵‵}(fgh)(fhg)
= (fkk)^{2} − (feg)(fge) + Σ^{‵‵}(fkh)(fhk);  (119)

where h, under Σ^{‵‵}, receives only n − 4 values, being distinct from each of the four unequal indices, e f g k.

[18.] By changing e to f in (95), and attending to the properties of the symbols (fgh), we obtain the expression

(fg) = Σ(fgh)(hff);  (120)

where f and g are unequal, and the summation Σ extends from h = 1 to h = n. The term for which h = f vanishes, and the formula (120) may be thus more fully written:

(fg) − (fge)(eff) = (fgg)(gff) + Σ^{‵}(fgh)(hff);  (121)

where the letters e f g denote again some three unequal indices, and the summation Σ^{‵} is performed as in the foregoing paragraph. But also, by (97) and (100),

(fg) − (fge)(eff) = (fee)(gee) + Σ^{‵}(feh)(ghe);  (122)

subtracting, therefore, (122) from (121), we eliminate the symbol (fg), and obtain the type

II. …  (fee)(gee) − (fgg)(gff) = Σ^{‵}{(fgh)(hff) − (feh)(ghe)};  (123)

which represents in general a system of n(n − 1)(n − 2) distinct and homogeneous equations of the second dimension, containing each 2(n − 2) terms, and derived by eliminations of the kind last mentioned, from the formulæ (95), (97), (100), in a manner agreeable to the analysis of paragraph [14.]. Indeed, it was shown in that paragraph, that the equation

(gf) = (fg), (100),

though known from earlier and simpler principles to be true, might be regarded as included in (95) and (97); but this need not prevent us from using that equation in combination with the others, whenever it may seem advantageous to do so: and other combinations of them may with its help be formed, which are occasionally convenient, or even sometimes necessary, although the general results of the elimination of the symbols (fg) are sufficiently represented by the recent type II., or by the formula (123). For example, a subordinate type, including only \frac{1}{2}n(n − 1)(n − 2) distinct equations, of 2(n − 2) terms each, between the symbols (fgh), may thus be formed, by subtracting (95) from (97), under the condition that e f g shall still denote some three unequal indices; namely,

0 = Σ{(feh)(ghe) − (geh)(fhe)};  (124)

or more fully, but at the same time with the suppression of a few parentheses, which do not appear to be at this stage essential to clearness,

(fge)(eff + egg) = Σ^{‵}(geh . fhe − feh . ghe) :  (125)

this last formula admitting also of being obtained from (122), by interchanging f and g, and subtracting. Again, a type which is in general still more subordinate, as including only \frac{1}{2}n(n − 1) distinct equations, of 2(n − 2) terms each, may be derived by the same process from (120); namely the type,

0 = Σ(fgh)(hff + hgg);  (126)

or in a slightly more expanded form,

(fge)(eff + egg) = Σ^{‵}(gfh)(hff + hgg);  (127)

which may also be easily derived, in the same way, from (121). It will, however, be found, by pursuing a little further the analysis of [14.], that the equations of this last type, (126) or (127), are always consequences of the equations of the intermediate type, (124) or (125); the sum of the n − 2 equations of the form (125), which answer to the various values of e that remain when f and g have been selected, being in fact equivalent to the formula (126). It will also be found, by the same kind of analysis, that the intermediate equations of the type (124) or (125) are generally deducible from those of the form (123). But on the subject of these general reductions, connected with the elimination of the symbols (fg) or (gf), it may be proper to add a few words.

[19.] Let us admit, at least as temporary abridgements, the notations

[fg] = Σ(fgh . hff); [fge] = Σ(fhe . geh);  (128)

where e, f, g are any three unequal indices, and h varies under Σ, as before, from 1 to n. Then the formula (95) gives n − 1 distinct equivalents for the symbol (fg), of which one is by (120) of the form [fg], and the n − 2 others are each of the form [fge]; in such a manner that we may write, instead of (95), with these last notations, the system of the two formulæ,

(fg) = [fg], (fg) = [fge];  (129)

whereof the latter is equivalent to a system of n − 2 equations: and of course, instead of (97), we may in like manner write

(gf) = [gf], (gf) = [gfe].  (130)

The equations (99) may now be thus presented:

\begin{matrix}(n − 1)(fg) & = [fg] + Σ^{‵}[fge] = ΣΣ(fhe . geh); \\ (n − 1)(gf) & = [gf] + Σ^{‵}[gfe] = ΣΣ(feh . ghe); \end{matrix} \}  (131)

where e under the sign Σ′ is distinct from each of the two indices f and g; but, under the double sign ΣΣ, both e and h may each receive any one of the values from 1 to n. The two double sums are equal, as in [14.], and therefore we must have, identically,

[fg] + Σ′[fge] = [gf] + Σ′[gfe];  (132)

the equation (100) being at the same time seen again to be a consequence, by simple additions, of the formulæ (95) and (97). Thus, after assigning any two unequal values to the indices f and g, we see that the two symbols, (fg), (gf); the two others, [fg], [gf]; the n − 2 symbols, [fge]; and the n − 2 symbols, [gfe], are indeed all equal to each other: but that the 2n − 1 equations between these 2n equal symbols are connected by a relation, such that any 2n − 2 of them, which are distinct among themselves, include the remaining one: and that therefore, after the elimination of (fg) and (gf), there remain only 2(n − 2) distinct equations of condition, as was otherwise shown in [14.]. But, in that paragraph, we proposed to form those resulting conditions on a plan which may now be represented by the formulæ

[fg] = [fge], [gf] = [gfe];  (133)

whereas we now prefer, for the sake of the convenience gained by the disappearance of certain terms in the subtractions, to employ that other mode of combination, which conducted in [18.] to the formula (123), and may now be denoted as follows:

[fg] = [gfe], [gf] = [fge].  (134)

Summing these last with respect to e, we find

(n − 2)[fg] = Σ′[gfe], (n − 2)[gf] = Σ′[fge];  (135)

and therefore, by the identity (132),

(n − 3)[gf] = (n − 3)[fg].  (136)

If, then, n > 3, we are entitled to infer, from (123) or (134), the following formula, which is equivalent to (126),

[gf] = [fg];  (137)

and therefore also by (134) this other type, equivalent to (124),

[fge] = [gfe],  (138)

which includes n − 2 equations, when f and g are given, and conducts, reciprocally, by (132), to (137). In general, therefore, if we adopt the type (134), we need not retain also either of these two latter types, (137), (138). But in the particular case where n = 3, that is, in the case of quadrinomes, the identity (132) reduces the two equations (134) to one, after f and g have been selected; and with this one we must then combine either of the two equations (137) or (138), which in this case become identical with each other.

[20.] In particular, for this case of quadrinomials (n = 3), we have with the notations (114), (128), the four following values for (23), or for b_{1} (compare again a note to [12.]):

\begin{matrix}[23] & = 231 . 122 + 233 . 322 = l_{1}n_{1} − n_{2}m_{3}; \\ [231] & = 211 . 311 + 231 . 313 = −m_{2}n_{3} + l_{1}m_{1}; \\ [32] & = 321 . 133 + 322 . 233 = l_{1}m_{1} − m_{3}n_{2}; \\ [321] & = 311 . 211 + 321 . 212 = −n_{3}m_{2} + l_{1}n_{1}; \end{matrix} \}  (139)

but, whether we equate the first to the fourth, or the second to the third of these expressions for b_{1}, in conformity with the type (134), we obtain only one common equation of condition, n_{2}m_{3} = n_{3}m_{2}, equivalent indeed by cyclical permutation to three, namely to the following,

0 = n_{2}m_{3} − n_{3}m_{2} = n_{3}m_{1} − n_{1}m_{3} = n_{1}m_{2} − n_{2}m_{1};  (140)

which evidently agree with certain simple combinations of the six equations on the two last lines of (86). If however we compare either the first value (139) with the third, or the second of those values with the fourth, according to the type (137) or (138), we find by each comparison the common condition l_{1}n_{1} = l_{1}m_{1}, and thus recover the equations of the first line of (85). In this way then we may obtain the required number of six distinct equations, with two terms each, between the nine symbols (fgh), or l_{1} … n_{3}, for the case of quadrinomes, by elimination of the three symbols (fg), or of b_{1}, b_{2}, b_{3}.

[21.] For the case of quines (n = 4), the general theory requires that the corresponding elimination of the 6 = \frac{1}{2}n(n − 1) symbols of this form (fg), or b_{1} … c_{3}, should conduct to 24 = n(n − 1)(n − 2) distinct equations of condition, with 4 = 2(n − 2) terms each, between the \frac{1}{2}n^{2}(n − 1) = 24 symbols of the form (fgh), or l_{1} … u_{3}, each equation thus obtained being homogeneous, and of the second dimension; and that all these 24 conditions should be included in the formula (134), or in the single type (123). And in fact we thus obtain, by comparison of the six expressions for b_{1}, of which one is

b_{1} = (23) = [23] = Σ(23h . h22) = l_{1}n_{1} − n_{2}m_{3} − p_{1}r_{2},  (141)

the four following equations of condition, included in that type or formula:

\begin{matrix}0 & = [23] − [321] = [32] − [231]; \\ 0 & = [23] − [324] = [32] − [234]; \end{matrix} \}  (142)

that is, with the notations l_{1} … u_{3},

\begin{matrix}n_{2}m_{3} − m_{2}n_{3} & = p_{3}s_{3} − p_{1}r_{2} = p_{1}r_{3} − p_{2}t_{2}; \\ n_{2}m_{3} + u_{2}u_{3} & = p_{3}s_{3} + l_{1}m_{1} = l_{1}n_{1} − p_{2}t_{2}; \end{matrix} \}  (143)

while we have in like manner six expressions for c_{1}, of which one is

c_{1} = [41] = Σ(41h . h44) = −(r_{1}u_{1} + s_{1}u_{2} + t_{1}u_{3}),  (144)

and of which the comparison conducts to the four other distinct conditions:

\begin{matrix}r_{1}u_{1} − n_{1}r_{2} & = l_{3}t_{3} + n_{3}t_{1} = −l_{2}s_{2} − t_{1}u_{3}; \\ r_{1}u_{1} + m_{1}r_{3} & = l_{3}t_{3} − s_{1}u_{2} = −l_{2}s_{2} − m_{2}s_{1}; \end{matrix} \}  (145)

where cyclical permutation of indices is still allowed. The equations obtained from the types (137), (138) would be found (as the theory requires) to be merely consequences of these; for example, by making e = 1, f = 2, g = 3, h = 4, those two types give only the conditions,

l_{1}(n_{1} − m_{1}) = p_{1}(r_{2} + r_{3}) = p_{2}t_{2} + p_{3}s_{3},  (146)

which are obviously included in (143).

[22.] With respect to those other homogeneous equations of the second dimension, between the symbols (fgh), which are obtained immediately, or without any elimination of the symbols (f), (fg), from the general conditions of association, and are included in the formula (96), they may now be developed as follows.

Making k = f in (96), and then interchanging f and e, for the sake of comparison with (123), we obtain the type

III. …  fgg . gee − gff . fee = Σ^{‵}(gfh . hee);  (147)

which includes generally \frac{1}{2}n(n − 1)(n − 2) distinct equations, of n − 1 terms each. For quines, we have thus 12 equations of 3 terms sufficiently represented by the following:

\begin{matrix}n_{1}n_{2} − m_{2}m_{1} & = p_{3}r_{3}, & n_{1}u_{2} + m_{2}u_{1} & = −l_{3}u_{3}; \\ r_{2}m_{2} + u_{2}r_{1} & = s_{2}n_{3}, & r_{2}n_{2} − u_{2}r_{3} & = +t_{2}m_{1}; \end{matrix} \}  (148)

the value 4 being attributed to the index h or e, in forming the equations on the first line, but to f or g for the second line. For quadrinomes, the corresponding equations are only three, namely

0 = n_{1}n_{2} − m_{1}m_{2} = n_{2}n_{3} − m_{2}m_{3} = n_{3}n_{1} − m_{3}m_{1};  (149)

which however are sufficient, in conjunction with the three lately marked as (140), to reproduce the six equations of the two last lines of (86). In general, by adding and subtracting the two types (123), (147), we obtain the formula,

(fee ± fgg)(gee ∓ gff) = Σ^{‵}(fgh)(hff ∓ hee) − Σ^{‵}(feh . ghe);  (150)

where, as a verification, if we take the lower signs, and interchange f and g, so as to recover the first member with the upper signs, the comparison of the two expressions for that member conducts to an equation between the two second members, which may also be obtained by the comparison of (125) and (127).

[23.] Again, making f = e in (96), and then changing k to f, we obtain the formula,

IV. …  0 = (egf)(eff + egg) + Σ^{‵}(egh . ehf);  (151)

where e f g h are again unequal indices. This IVth type includes generally n(n − 1)(n − 2) distinct equations, with n − 1 terms each. For the case n = 3 there arise thus 6 equations of 2 terms, namely the six on the two last lines of (85); so that the 18 equations (85), (86), for associative quadrinomials, have thus been completely reproduced, as consequences of the general theory. For the case of quines, the type (151) gives 24 equations of 3 terms, which may be represented as follows:

\begin{matrix}0 & = l_{2}(n_{1} − m_{1}) + p_{2}s_{1} = l_{3}(n_{1} − m_{1}) + p_{3}t_{1}; \\ 0 & = s_{2}(r_{2} + r_{3}) + t_{1}t_{2} = t_{3}(r_{2} + r_{3}) + s_{3}s_{1}; \end{matrix} \}  (152)

\begin{matrix}0 & = t_{1}(u_{1} − m_{1}) + l_{3}s_{1} = p_{2}(u_{1} − m_{1}) + l_{2}p_{3}; \\ 0 & = s_{1}(u_{1} + n_{1}) − l_{2}t_{1} = p_{3}(u_{1} + n_{1}) − l_{3}p_{2}; \end{matrix} \}  (153)

either h or e being = 4 in (152), and either f or g having that value in (153), while 1, 2, 3 may still be cyclically permuted.

[24.] Finally, by supposing, as in (119), that e f g k are four unequal indices, and that h under Σ^{‵‵} is unequal to each of them, we obtain from (96) one other type, including generally \frac{1}{2}n(n − 1)(n − 2)(n − 3) equations, of n − 1 terms each, but furnishing no new conditions of association for quadrinomials: namely,

V. …  efk . gee + fgk . egg + egk . fkk = Σ^{‵‵}(fhk . geh).  (154)

For quines, the sum Σ^{‵‵} vanishes, and we obtain twelve equations of three terms each, which may (with the help of permutations) be all represented by the four following:

\begin{matrix}0 & = l_{1}r_{2} − u_{2}s_{3} − n_{3}t_{2} = l_{1}r_{3} − m_{2}s_{3} + u_{3}t_{2}, \\ 0 & = l_{1}r_{1} + n_{2}s_{3} + m_{3}t_{2} = n_{1}p_{1} − m_{2}p_{2} − u_{3}p_{3}; \end{matrix} \}  (155)

where the index 4 has been made to coincide with e or with g in the first line, but with f or k in the second.

[25.] In general, the number of distinct associative equations, included in the three last types (147), (151), (154), or III., IV., V., which have been all derived from the formula (96), and have been obtained without elimination of (f) or (fg), amounts in the aggregate to

\frac{1}{2}n(n − 1)(n − 2) + n(n − 1)(n − 2) + \frac{1}{2}n(n − 1)(n − 2)(n − 3) = \frac{1}{2}n^{2}(n − 1)(n − 2); (156)

as, by the analysis of [14.], it ought to do. And when we add this number to the n(n − 2) of the type I., or (113), and to the n(n − 1)(n − 2) of the type II., or (123), obtained by such elimination, we have in all this other number,

\frac{1}{2}n^{2}(n − 1)(n − 2) + n^{2}(n − 2) = \frac{1}{2}n^{2}(n + 1)(n − 2),  (157)

of distinct and homogeneous equations of the second dimension, between the \frac{1}{2}n^{2}(n − 1) symbols of the form (fgh): as, by the formulæ (109), (110) of [15.], we ought to have. As regards the signification of the five foregoing principal types, which it has been thought convenient to distinguish among themselves, and to arrange according to the various ways in which they involve the symbols of the form (eff), it will be found, on reviewing the analysis employed, that they all express ultimately only consequences of that one very simple and useful formula,

Vι Vι′ι′′ = ι′′ Sιι′ − ι′ Sι′′ι, (67)

which, with a slightly different notation, has been elsewhere shown to be so important in the Calculus of Quaternions*. In fact, the equations (95) … (98), on which those five separate types have been founded, may all be deduced from (67) and (69), whereof the latter is a consequence of the former.

Observatory of Trinity College, Dublin, August 14, 1854.

* See the author’s Lectures on this subject, cited in former notes to this paper.

Section IV.

[26.] For quines, the equations of condition between the 24 symbols l_{1} … u_{3} amount (as has already been remarked) to 80 in all; namely to 8, 24, 12, 24, and 12 equations, included respectively in the five types last mentioned, and sufficiently developed above, by the formulæ (117) (118) (143) (145) (148) (152) (153) (155): which also enable us, with the help of (141), (144), to determine the values of the four symbols a_{1} … a_{4}, and of the six other symbols b_{1} … c_{3}, when values of l_{1} … u_{3} have been found, which satisfy the eighty conditions. And then, if we denote the quine itself by the following expression (compare [1.]),

P = x_{0} + ι_{1}x_{1} + ι_{2}x_{2} + ι_{3}x_{3} + ι_{4}x_{4},  (158)

which is a little more symmetric than the form (111), the laws of multiplication of any two such quines, P, P′, will be sufficiently expressed by the formulæ

\begin{matrix}ι_{1}^{2} = a_{1}, ι_{4}^{2} = a_{4}, Sι_{2}ι_{3} = b_{1}, Sι_{1}ι_{4} = c_{1}, \\ Vι_{2}ι_{3} = ι_{1}l_{1} + ι_{2}m_{3} + ι_{3}n_{2} + ι_{4}p_{1}, \\ Vι_{1}ι_{4} = ι_{1}r_{1} + ι_{2}s_{1} + ι_{3}t_{1} + ι_{4}u_{1}; \end{matrix} \}  (159)

if we remember that 1, 2, 3 may still be cyclically permuted, and that the law of conjugation (32) gives

Kι′ι = ιι′, Sι′ι = Sιι′, Vι′ι = −Vιι′.  (160)

For in this manner, by (41), if ϖ denote, as in (14), the vector part of P, so that

ϖ = ι_{1}x_{1} + ι_{2}x_{2} + ι_{3}x_{3} + ι_{4}x_{4},  (161)

we shall have

Sϖϖ′ = a_{1}x_{1}x′_{1} + a_{2}x_{2}x′_{2} + a_{3}x_{3}x′_{3} + a_{4}x_{4}x′_{4}
+ b_{1}(x_{2}x′_{3} + x_{3}x′_{2}) + &c. + c_{1}(x_{1}x′_{4} + x_{4}x′_{1}) + &c.,  (162)
Vϖϖ′ = (ι_{1}l_{1} + ι_{2}m_{3} + ι_{3}n_{2} + ι_{4}p_{1})(x_{2}x′_{3} − x_{3}x′_{2}) + &c.
+ (ι_{1}r_{1} + ι_{2}s_{1} + ι_{3}t_{1} + ι_{4}u_{1})(x_{1}x′_{4} − x_{4}x′_{1}) + &c.,  (163)

each “&c.” representing terms obtained by the permutations already mentioned; and if the constants a b c l m n p r s t u have been chosen so as to fulfil the conditions above developed, we may then conclude (compare (51)) that the following equations of association hold good, for the multiplication of any three such vector-units ι, or quadrinomial vectors ϖ, or quinquinomial expressions P, whether equal or unequal among themselves:

ι . ι′ι′′ = ιι′ . ι′′; ϖ . ϖ′ϖ′′ = ϖϖ′ . ϖ′′; P . P′P′′ = PP′ . P′′;  (164)

which it has been the main object of our recent investigations to establish.

[27.] Without pretending to do more, on the present occasion, than merely to exemplify the possibility of satisfying, for quines, the foregoing equations of association, I may here remark that if we restrict the question by assuming (with the usual permutations),

(A, B)* …  n_{1} = m_{1}, p_{1} = 0, u_{1} = 0,  (165)

then numerous simplifications take place, and the 80 equations between the 24 symbols l m n p r s t u are found to reduce themselves to 44 equations between the 15 symbols l m r s t, obtained from the five types I. to V. of recent paragraphs, which may be thus denoted and arranged:

from type I.  m_{1}^{2} = l_{2}l_{3}, r_{2}^{2} − r_{1}^{2} = s_{3}t_{1} − s_{2}t_{3};  (166)
from II. and III.  m_{1}l_{1} = m_{2}m_{3}, m_{1}r_{1} = m_{2}s_{1} = m_{3}t_{1},  (167)
and  m_{1}r_{2} = l_{2}s_{2}, m_{1}r_{3} = l_{3}t_{3}, m_{1}(r_{1} + r_{2} + r_{3}) = 0; (168)
from IV.  s_{2}(r_{2} + r_{3}) = −t_{1}t_{2}, t_{3}(r_{2} + r_{3}) = −s_{3}s_{1},  (169)
and  m_{1}s_{1} = l_{2}t_{1}, m_{1}t_{1} = l_{3}s_{1};  (170)
and from V.  m_{1}s_{2} = l_{3}r_{2}, m_{1}t_{3} = l_{2}r_{3}, l_{1}(r_{1} + r_{2} + r_{3}) = 0. (171)

Now these conditions may all be satisfied in each of two principal ways, conducting to two distinct systems of associative quines, which may be called Systems (A) and (B), but which are not the only possible systems of such quines, because we need not have commenced by assuming the equations (165), although that assumption has simplified the problem. For first we may suppose that the constants l and m are different from zero, but that the constants r are connected by the relation

(A) …  r_{1} + r_{2} + r_{3} = 0;  (172)

or secondly, we may reject this relation between the constants r, and suppose instead that the six constants l and m all vanish, so that

(B) …  l_{1} = l_{2} = l_{3} = m_{1} = m_{2} = m_{3} = 0.  (173)

With the first supposition, (172), we are to combine the nine relations between the fifteen constants l m r s t, which are sufficiently expressed by the formula (167), or by the following:

(A_{1}) …  l_{1} = m_{1}^{−1}m_{2}m_{3}, s_{1} = m_{2}^{−1}m_{1}r_{1}, t_{1} = m_{3}^{−1}m_{1}r_{1};  (174)

and then all the other conditions of association will be found to be satisfied, if we equate each of the ten symbols a b c to zero, or if we establish this other formula,

(A_{2}) …  a_{1} = 0, b_{1} = 0, c_{1} = 0, a_{4} = 0 :  (175)

* This line is lettered thus, because it contains the conditions common to the two systems (A) and (B) of associative quines, which are deduced a little further on. while there will still remain five arbitrary constants of the system, for instance r_{1} r_{2} m_{1} m_{2} m_{3}. With the second supposition, (173), we are to combine four distinct relations between the nine constants r s t, contained in the formula (169), or in the following:

(B_{1})* …  r_{1} + r_{2} = −s_{1}^{−1}t_{1}t_{3}, s_{1}s_{2}s_{3} = t_{1}t_{2}t_{3};  (176)

which give also, as a consequence, this other relation:

(r_{1} + r_{2})(r_{2} + r_{3})(r_{3} + r_{1}) = −s_{1}s_{2}s_{3};  (177)

and then the other conditions of association will all be satisfied, if we make, instead of (175),

(B_{2}) …  a_{1} = b_{1} = c_{1} = 0; a_{4} = (r_{1} + r_{2} + r_{3})^{2} :  (178)

this system also involving five arbitrary constants, for example s_{1} s_{2} t_{1} t_{2} t_{3}. The assertion respecting quines, which was made near the end of [16.], has therefore been fully justified.

[28.] Finally, as regards the system (A) of quines, it may be observed, 1st, that in this system, by (162) and (175), we have generally,

(A_{3}) …  Sϖϖ′ = 0;  (179)

or that “the product of any two quadrinomial vectors ϖ, ϖ′, reduces itself to a pure vector;” and 2nd, that, by (163), (165), “this vector product, ϖϖ′, is of trinomial form, involving no part with ι_{4} for a factor.” This product is therefore already seen to be of the form

ϖϖ′ = ι_{1}X_{1} + ι_{2}X_{2} + ι_{3}X_{3};  (180)

but I say, 3rd, that “its three coefficients, or coordinates, X_{1}, X_{2}, X_{3} have constant ratios,” or that “the product ϖϖ′ may be constructed by a right line in space of which the direction though not the length is fixed,” and which may therefore be conceived to “coincide in position with one fixed axis (ξ) of the system.” In fact, by (163), (165), (174), we have

m_{1}X_{1} = m_{2}X_{2} = m_{3}X_{3} = X,  (181)

and therefore

(A_{4}) …  ϖϖ′ = Xξ,  (182)

if we make for abridgement

X = m_{2}m_{3}(x_{2}x′_{3} − x_{3}x′_{2}) + m_{3}m_{1}(x_{3}x′_{1} − x_{1}x′_{3}) + m_{1}m_{2}(x_{1}x′_{2} − x_{2}x′_{1})
(A_{5}) … { + m_{1}r_{1}(x_{1}x′_{4} − x_{4}x′_{1}) + m_{2}r_{2}(x_{2}x′_{4} − x_{4}x′_{2}) + m_{3}r_{3}(x_{3}x′_{4} − x_{4}x′_{3}), (183)
and ξ = m_{1}^{−1}ι_{1} + m_{2}^{−1}ι_{2} + m_{3}^{−1}ι_{3}.  (184)

* It must be observed that these equations (176), which are part of the basis of the system (B), are true in the system (A) also, as corollaries from (174) and (172) which last equation does not hold in (B); and which allows us to reduce, for (A) but not for (B), the relation (177) to the simpler form r_{1}r_{2}r_{3} = s_{1}s_{2}s_{3}.

In the 4th place, “if any quadrinomial vector ϖ be multiplied by or into the axis ξ, the product vanishes;” or in symbols,

(A_{6}) …  ξϖ = 0, ϖξ = 0;  (185)

because by (172) the scalar coefficient X becomes = 0, if we change either x_{1}, x_{2}, x_{3}, and x_{4}, or x′_{1}, x′_{2}, x′_{3}, and x′_{4}, to m_{1}^{−1}, m_{2}^{−1}, m_{3}^{−1}, and 0, respectively. This coefficient X vanishes also, when we equate x′_{1} x′_{2} x′_{3} x′_{4} to x_{1} x_{2} x_{3} x_{4} respectively; and hence, or from (179), we may infer, 5th, that “in this system of quines (A), the square of every quadrinomial vector vanishes.” And finally, by an easy combination of the formulæ (182), (185), or of the 3rd and 4th of the foregoing properties of this system, we see, 6th, that, in it, “every product of three quadrinomial vectors vanishes;” or that

(A_{7}) …  ϖϖ′ . ϖ′′ = 0, ϖ . ϖ′ϖ′′ = 0.  (186)

[29.] The associative property (164) is therefore verified for the system (A), by showing that, in it, each of the two ternary products of vectors, which ought to be equal, vanishes. In the system (B), it is easy to see that any such ternary product must be itself a vector; because, in (B), no binary product of vectors involves ι_{4}, nor does any such product involve a scalar part, except what arises from ι_{4}^{2}. We have, therefore, here, this new result,

(B_{3}) …  S(ϖϖ′ . ϖ′′) = S(ϖ . ϖ′ϖ′′) = 0.  (187)

And when we proceed to develop these two ternary products, the associative property of multiplication is again found to be verified, under the form,

(B_{4}) …  ϖϖ′ . ϖ′′ = ϖ . ϖ′ϖ′′ = a_{4}(ϖx′_{4}x′′_{4} − ϖ′x′′_{4}x_{4} + ϖ′′x_{4}x′_{4});  (188)

where it is worth observing that, by the laws of the system in question, the result may be put under this other and somewhat simpler form:

(B_{5}) …  ϖϖ′ . ϖ′′ = ϖ . ϖ′ϖ′′ = ϖ Sϖ′ϖ′′ − ϖ′ Sϖ′′ϖ + ϖ′′ Sϖϖ′.  (189)

Indeed, this last expression might have been foreseen, as a consequence from the general principles of this whole theory of associative polynomes*, combined with the particular property (187) of the quines (B). For, by that property, each of the two ternary products is equal to its own vector part; but by (101) we have, generally, in the present theory, as in the calculus of quaternions, the following expression for the vector part of the product of any three vectors, of any such associative polynomes as we are considering:

V . ρστ = ρ Sστ − σ Sτρ + τ Sρσ;  (190)

which is a formula of continual application in Quaternions, and in these extensions also is important.

Observatory of Trinity College, Dublin, August 16, 1854.

* It will hereafter be proved generally that for all associative polynomes which satisfy the law of conjugation (though not exclusively for such associative polynomes), the tensor, as defined in [6.], is also a modulus; which theorem can be verified for the quines (A) and (B), and for the quadrinomes and tetrads so lettered in [13.], as well as for the trinomes [11.].

Section V.

[30.] In applying to associative quines the general theory of the Third Section*, we may (as has been seen) omit each of the signs Σ^{‵} as unnecessary, the index h receiving only one value in the sum thereby indicated; and may suppress each sum Σ^{‵‵} as vanishing. In this manner the type IV., or the formula (151), becomes,

IV. …  (egf)(eff + egg) = geh . ehf;  (191)

while the equation (127), already derived as a sub-type from II., gives, by interchanging e and f,

(egf)(fee + fgg) = (geh)(hee + hgg).  (192)

Multiplying the latter of these two equations by eff + egg, and the former by fee + fgg, and subtracting, we eliminate the symbol (egf), and find that

(geh){(ehf)(fee + fgg) − (eff + egg)(hee + hgg)} = 0;  (193)

and a similar elimination of (geh) gives the equation,

(egf){(ehf)(fee + fgg) − (eff + egg)(hee + hgg)} = 0.  (194)

And because (geh) = −(egh), by (34), we may make any separate or combined interchanges, of e with g, and of f with h, and so vary the expression within the { }, without introducing any new factor, distinct from (egf) and (egh), outside them. If, then, for any particular arrangement of the four unequal indices, e, f, g, h, as chosen from among the four numbers 1, 2, 3, 4, the two following conditions are not both satisfied,

(egf) = 0, (egh) = 0,  (195)

we must have, for that arrangement of the indices, a system of four other equations, whereof one is

VI. …  (ehf)(fee + fgg) = (eff + egg)(hee + hgg),  (196)

while the three others are formed from it, by the interchanges just now mentioned. And if we further suppose that the two sums, fee + fgg and hee + hgg, are for the same arrangement different from zero, and write for abridgement, as a definition,

(ehf)_{0} = (fee + fgg)^{−1}(eff + egg)(hee + hgg),  (197)

the four equations furnished by the formula (196), which may be regarded as a sixth type for quines, may be concisely expressed as follows:

\begin{matrix}(ehf) & = (ehf)_{0}, & (ghf) & = (ghf)_{0}; \\ (efh) & = (efh)_{0}, & (gfh) & = (gfh)_{0}. \end{matrix} \}  (198)

* Comprising paragraphs [17.] to [25.], and published in the Phil. Mag. for October, 1854.

With the notations l_{1} … u_{3}, for the symbols (efg), (eff), we find thus that unless l_{1} and p_{1} both vanish, we must have the four equations,

\begin{matrix}t_{2}(n_{1} − m_{1}) & = (m_{2} − n_{2})(r_{2} + r_{3}); & s_{3}(n_{1} − m_{1}) & = (m_{3} − n_{3})(r_{2} + r_{3}); \\ p_{3}(r_{2} + r_{3}) & = (u_{2} + n_{2})(n_{1} − m_{1}); & p_{2}(r_{2} + r_{3}) & = (m_{3} − u_{3})(n_{1} − m_{1}); \end{matrix} \}  (199)

and that unless s_{1} = 0, t_{1} = 0, then

\begin{matrix}−l_{2}(u_{2} − m_{2}) & = (n_{1} + u_{1})(n_{3} + u_{3}); & t_{3}(u_{2} − m_{2}) & = (r_{1} + r_{2})(n_{3} + u_{3}); \\ l_{3}(n_{3} + u_{3}) & = (u_{1} − m_{1})(u_{2} − m_{2}); & s_{2}(n_{3} + u_{3}) & = (r_{1} + r_{3})(u_{2} − m_{2}). \end{matrix} \}  (200)

[31.] Supposing then that no one of the twelve symbols (efg) vanishes, and that each of the twelve sums eff + egg is also different from zero, the various arrangements of the four indices e f g h give us a system of twenty-four equations, included in the new type VI., or in any one of the four formulæ (198); which equations may, by (34), be arranged in twelve pairs, as follows:

(ehf) = (ehf)_{0} = −(hef)_{0}.  (201)

It might seem that twelve equations between the twelve symbols of the form (eff) should thus arise, by the comparison of two expressions for each of the twelve symbols of the form (efg); but if we write for abridgement

[g] = (fee + fgg)(fhh + fgg){(ehf)_{0} + (hef)_{0}},  (202)

and observe that by the definition (197) of the symbol (ehf)_{0}, we have then

[g] = (eff + egg)(fhh + fgg)(hee + hgg) + (hff + hgg)(fee + fgg)(ehh + egg), (203)

we shall see that this quantity [g] is independent of the arrangement of the three indices e, f, h; and therefore that the twelve equations between the twelve symbols (eff), obtained through (201), reduce themselves to the four following relations,

[e] = 0, [f] = 0, [g] = 0, [h] = 0;  (204)

which are not even all distinct among themselves, since any three of them include the fourth. An easy combination of the two first or of the two last of these four relations (204) conducts to this other formula, which is equivalent to three distinct equations:

(eff + ehh)(fee + fgg)(ghh + gee)(hgg + hff)
= (eff + egg)(fee + fhh)(ghh + gff)(hgg + hee);  (205)

and which may also be thus written,

(hef)_{0}(egf)_{0} = (ehf)_{0}(gef)_{0}.  (206)

With the notations l_{1} … u_{3}, the twenty-four equations (201) are sufficiently represented by the formulæ (199) and (200), if cyclical permutation of the indices be employed; the four equations (204) take the forms,

\begin{matrix}(r_{1} + r_{2})(n_{3} + u_{3})(m_{2} − n_{2}) & = (r_{1} + r_{3})(m_{2} − u_{2})(m_{3} − n_{3}), \\ (n_{1} + u_{1})(n_{2} + u_{2})(n_{3} + u_{3}) & = (m_{1} − u_{1})(m_{2} − u_{2})(m_{3} − u_{3}); \end{matrix} \}  (207)

whereof the equation on the second line may be obtained from the product of the three represented by the first line: and the three equations (205) or (206) are included in the following, which is evidently a consequence of (207),

(r_{1} + r_{3})(n_{1} + u_{1})(n_{2} + u_{2})(n_{3} − m_{3}) = (r_{1} + r_{2})(u_{1} − m_{1})(n_{2} − m_{2})(u_{3} − m_{3}). (208)

[32.] As regards the quotients and products of the symbols (efg), which we shall continue to write occasionally without parentheses, we have by type VI., or by (197), (198),

\frac{ehf}{ghf} = \frac{eff + egg}{gff + gee};  (209)
ehf . efh = (eff + egg)(ehh + egg);  (210)
ehf . gfh = (eff + egg)(ghh + gee);  (211)

eliminating (ehf) between the two last of which three equations, we obtain a relation of the same form as the first. Interchanging g and h in (210), and subtracting, we find that

I. …  ehf . efh − egf . efg = (egg)^{2} − (ehh)^{2};  (212)

but this is precisely what the type I., or the formula (113), becomes for quines, when we cyclically advance the four indices in the order f e g h; the conditions (177), (118) of that first type will therefore be satisfied, if we satisfy those of the sixth. Had we divided instead of subtracting, we should have found

\frac{ehf . efh}{egf . efg} = \frac{eff + egg}{eff + ehh}.  (213)

To interchange f, g, and divide, would only lead by (210) to another equation of this last form; but the same operations performed on (211) conduct to the equation

\frac{ehf}{heg} = \frac{ghh + gee}{fhh + fee};  (214)

which, when we interchange g and h, reproduces the formula (192); and shows thereby that the sub-type (127), included under type II., is satisfied by our new type VI., which indeed it had assisted to discover. The same equation (192) may also be derived from the formula (205), by dividing each member of that formula by (fee + fhh)(hff + hgg), and attending to the expressions given by type VI., for (egf) and (geh) respectively. To interchange e h, in (211) and divide, would only conduct to another equation of the same form as (214). Permuting cyclically the three indices e, f, g in (209), and multiplying together the two equations so obtained therefrom, the product gives

\frac{fhg . ghe}{ehg . fhe} = \frac{gff + gee}{eff + egg};  (215)

and if we multiply this equation by (209) itself, we find that

ehf . fhg . ghe = ghf . ehg . fhe.  (216)

In fact if we operate thus on the expression (197) for (efh)_{0}, or for its equal (ehf), or on the formula (196), we are led to this new equation,

ehf . fhg . ghe = (hee + hgg)(hff + hee)(hgg + hff),  (217)

of which the second member does not alter, when we interchange any two of the three indices e, f, g. Another multiplication of three equations of the form (209), with the cycle e g h, conducts to the equation [f] = 0 of (204). Interchanging e, h in (210), and substituting the value so obtained for the product of the two extreme factors of the second member of (217), we find this other expression,

ehf . fhg . ghe = hef . hfe . (hee + hff);  (218)

which is still seen to remain unaltered, by an interchange of e and f. Interchanging f, g, and dividing, we obtain by (216) an equation of the same form as (213); and if we divide each member of (218) by (hef), we are conducted to the formula

IV. …  fhg . hge = hfe . (hee + hff),  (219)

which is of the same form as the equation (191), or as the type IV., and may be changed thereto by cyclical permutation of the four indices e f g h. The same relation (219) may also be derived more directly from type VI., by substitutions of the values (198); for it will be found that the definition (197) gives this identity,

(fhg)_{0}(ghe)_{0} = (hee + hff)(fhe)_{0}.  (220)

The conditions of type IV., like those of type I., and of the subtype (127) of II., are therefore all included in those of the new type VI.; which gives also in various ways this other formula respecting products of four symbols of the form (efg),

egh . fhg . gfe . hef = ehg . fgh . gef . hfe :  (221)

indeed it will be found that the members of this last equation, taken in their order, are respectively equal by (196) to the members of the equation (205).

With the notations l_{1} … u_{3}, supposing that none of the twelve constants l p s t vanish, and that the twelve combinations of the forms n_{1} − m_{1}, n_{1} + u_{1}, u_{1} − m_{1}, r_{1} + r_{2}, are in like manner different from zero, we find thus, or from the equations (199) (200), combined with their consequences (207), the following among other relations, in which cylical permutation of the indices is still allowed:

\begin{matrix}\frac{l_{1}}{t_{2}} & = −\frac{n_{3} + u_{3}}{r_{1} + r_{3}} , & \frac{t_{2}}{s_{3}} & = \frac{m_{2} − n_{2}}{m_{3} − n_{3}} , & \frac{s_{3}}{l_{1}} & = \frac{r_{1} + r_{2}}{u_{2} − m_{2}} , \\ \frac{p_{1}}{p_{2}} & = \frac{m_{2} − u_{2}}{n_{1} + u_{1}} , & \frac{l_{1}}{p_{1}} & = \frac{r_{2} + r_{3}}{n_{1} − m_{1}} , & \frac{s_{1}}{t_{1}} & = \frac{n_{3} + u_{3}}{m_{2} − u_{2}} ; \end{matrix} \}  (222)

\begin{matrix}l_{2}l_{3} & = (n_{1} + u_{1})(m_{1} − u_{1}), & s_{2}t_{3} & = (r_{1} + r_{2})(r_{1} + r_{3}), \\ p_{1}s_{2} & = (m_{2} − n_{2})(m_{2} − u_{2}), & p_{1}t_{3} & = (m_{3} − n_{3})(n_{3} + u_{3}), \\ −l_{2}s_{2} & = (n_{1} + u_{1})(r_{1} + r_{3}), & l_{3}t_{3} & = (u_{1} − m_{1})(r_{1} + r_{2}), \\ p_{1}s_{1} & = (m_{1} − n_{1})(n_{3} + u_{3}), & p_{1}t_{1} & = (m_{1} − n_{1})(m_{2} − u_{2}). \end{matrix} \}  (223)

The conditions (152), (153), of the fourth type, are satisfied; and we have these other products, of which some have occurred already, in (176), (177), in connexion with the particular systems (A) and (B) of quines:

\begin{matrix}s_{1}s_{2}s_{3} = t_{1}t_{2}t_{3} & = −(r_{1} + r_{2})(r_{2} + r_{3})(r_{3} + r_{1}); \\ s_{1}p_{2}l_{3} = t_{1}l_{2}p_{3} & = (n_{1} + u_{1})(m_{1} − u_{1})(m_{1} − n_{1}); \\ s_{2}s_{3}l_{2}p_{3} & = t_{2}t_{3}p_{2}l_{3}; \end{matrix} \}  (224)

where the two members of the equation on the last line are easily proved by (223) to be respectively equal to those of (208).

[33.] As yet we have only partially satisfied the conditions of type II., or of the formula (123), which may for quines be written thus:

II. …  fgh . hff = fee . gee − fgg . gff + feh . ghe.  (225)

Substituting for the last product in this formula its value given by (211), namely

feh . ghe = (fgg + fhh)(gee + gff),  (226)

and writing for abridgement

v_{f} = fee + fgg + fhh,  (227)

we are in this way led to establish the following seventh type* for quines:

VII. …  fgh . hff = v_{f} . gee + fhh . gff.  (228)

* It will be shown that this single type (the seventh) includes all the others, or is sufficient to express all the general conditions of association, between the 24 symbols of the forms (efg) and (eff). But the eliminations required for this deduction cannot be conveniently described at this stage.

Or since, by the sixth type, we have already

VI. …  fgh . (hee + hff) = (fee + fhh)(gee + gff),  (229)

it is only necessary, for the purpose of satisfying the conditions of type II., or the equations (143), (145), to suppose besides that

III. …  fgh . hee = fee . gff − fgg . gee;  (230)

such being the expression which remains, when we subtract (228) from (229). But this last equation (230) is precisely what the type III., or the formula (147), becomes for quines; it reproduces therefore the equations (148): and conversely, if we retain that old type III., it will not be necessary, although it may be convenient, to introduce the new type VII., in combination with type VI. And if in (230) we substitute for the symbol (fgh) its value given by (229), and so combine types III. and VI., we obtain the equation

\frac{fee + fhh}{hee + hff} = \frac{gff . (fee + fgg) − fgg . (gee + gff)}{hee . (gee + gff)};  (231)

that is, by (209),

\frac{fge}{hge} + \frac{fgg}{hee} = \frac{gff}{hee} . \frac{fhe}{ghe};  (232)

or finally,

V. …  hee . fge = fgg . ghe − gff . fhe.  (233)

But this is exactly what the type V., or the formula (154), becomes for quines, when we suppress the sum Σ^{‵‵}, change k to h, and advance cyclically the three indices e f h; it includes therefore the equations (155), which were the only remaining conditions of association to be fulfilled. If then we satisfy the two types, III. and VI., we shall satisfy all the conditions of association for quines: since we shall thereby have satisfied also the four other earlier types, namely those numbered as I. II. IV. V. It only remains then to consider what new restrictions on the constants (eff) are introduced by the comparison of the values which type III. gives for the other constants (efg), as expressed in terms of them, with the values furnished by type VI.; or to discuss the consequences of the following general formula, obtained by eliminating the symbol (fgh) between (229) (230), and not essentially distinct from the recent equations (231):

VIII. …  hee . (fee + fhh)(gee + gff) = (hee + hff)(fee . gff − gee . fgg); (234)

which contains all the old and new relations, subsisting between the twelve constants of the form (eff), and may be regarded as an eighth type for quines.

[34.] Denoting the first minus the second member of (234) by the symbol [efgh], we easily see that

[efgh] = gee . (v_{f} . hee + fgg . hff) + gff . (fhh . hee − hff . fee)
= hee . (v_{f} . gee + fhh . gff) + hff . (fgg . gee − gff . fee);  (235)

and therefore that we have, identically,

[efgh] = [efhg];  (236)

this last or eighth type (234) contains therefore, at most, only a system of twelve equations. Interchanging f and g, and attending to the notation (202), we see, by (203), (234), that of the three equations

[efgh] = 0, [egfh] = 0, [e] = 0,  (237)

any two include the third; if then we only seek what new conditions, additional to those marked (204), are to be satisfied by the symbols (eff), or rather by the eight following ratios of those symbols,

eff : egg : ehh; fee : fgg : fhh; gee : gff : ghh; hee : hff : hgg,  (238)

we need only retain at most four new equations, suitably selected from among those furnished by type VIII., such as the four following, which differ among themselves by the initial letters within the brackets, and so belong to different groups,

[efgh] = 0, [fghe] = 0, [ghef] = 0, [hefg] = 0;  (239)

and then to combine these with any three of the four former relations (204), for example with the three first, namely

[e] = 0, [f] = 0, [g] = 0;  (240)

from which the fourth equation [h] = 0 would follow, by means of the identity,

(ehf)_{0}(fge)_{0}(gfh)_{0}(heg)_{0} = (ehg)_{0}(fgh)_{0}(gfe)_{0}(hef)_{0}
= (eff + egg)(fee + fhh)(gee + ghh)(hff + hgg). (241)

It might seem however that the seven equations (239) and (240), thus remaining, should suffice to determine seven of the eight ratios (238): whereas I have found that it is allowed to assume two pairs of ratios arbitrarily, out of the four pairs (238), and then to deduce the two other pairs from them. For I find that it is sufficient to retain, instead of the twelve equations included under type VIII., or the seven equations (239), (240), a system of only four equations of the type just mentioned; namely two pairs, selected from any two of the four groups, which have (for each group, and also for each pair) a common initial letter within the brackets; for instance, these two pairs of equations:

[efgh] = 0, [egfh] = 0, [fegh] = 0, [fgeh] = 0;  (242)

which leave as many as eight arbitrary constants (for example these eight, eff, egg, ehh, fhh, ghh, hee, hff, hgg, from which all the rest can be determined) in the resulting system of associative quines. An outline of the investigation by which this important reduction is effected, may be presented in the following way.

[35.] The two first equations (242) connect the three last pairs of ratios (238), in such a manner that when any two of those three pairs are assumed, or known, the third can be determined. Hence, with the interpretation (197) of the symbol (ghf)_{0}, we easily find that those two equations (242) give,

(e) …  \{ \begin{matrix}gee . hgg − ghh . hee & = (ghf)_{0} . fee; \\ v_{g} . hee + gff . hgg & = (ghf)_{0} . fgg; \\ −(v_{h} . gee + hff . ghh) & = (ghf)_{0} . fhh; \end{matrix} \}  (243)

because we find that fee, fgg, ghh are proportional to the left-hand members of these last equations (243); and that the sum of the two first of those left-hand members is identically equal to the product (gee + gff)(hee + hgg). For the same reason, the two first of these three equations (243) express really only one relation, namely that which is contained in the second equation (242), although they do so under different forms, both of which it is useful to know; and it is convenient to have ready also this other combination, obtained by adding the three equations (243) together,

(e) …  v_{h} . gff − v_{g} . hff = v_{f} . (ghf)_{0};  (244)

which, like those equations (243) themselves, we shall consider as belonging to the group (e), because they are all derived from two of the three equations of that group, included under type VIII., which in the recent notation [efgh] have e for their initial letter; and because the third equation of that group, included under the same type, namely

(e) …  [ehfg] = 0,  (245)

may be derived from them, by the elimination of the symbol (ghf)_{0} between the first and third of the equations (243). In like manner the two last equations (242) include a third of the same type VIII., and belonging to the same group (f), namely

(f) …  [fheg] = 0;  (246)

because they conduct to the following system of expressions, which may be formed from (243)

(244) by cyclical permutation of efg:

(f) …  \{ \begin{matrix}eff . hee − ehh . hff & = gff . (ehg)_{0}; \\ v_{e} . hff + egg . hee & = gee . (ehg)_{0}; \\ −(v_{h} . eff + hgg . ehh) & = ghh . (ehg)_{0}; \\ v_{h} . egg − v_{e} . hgg & = v_{g} . (ehg)_{0}. \end{matrix} \}  (247)

Multiplying then the equations (243) and (244) by (ehg)_{0}, and observing that the identity (220) gives

(ehg)_{0} . (ghf)_{0} = (hee + hff)(ehf)_{0},  (248)

we find, on substitution of the first for the second members of (247), that the results are divisible by hee+hff; and that thus the elimination of the third pair of ratios (238), between (243), (244), (247), or between the four equations (242), conducts to expressions of the recent forms, namely,

(g) …  \{ \begin{matrix}hee . egg − ehh . hgg & = fgg . (ehf)_{0}; \\ v_{e} . hgg + eff . hee & = fee . (ehf)_{0}; \\ −(v_{h} . egg + hff . ehh) & = fhh . (ehf)_{0}; \\ v_{h} . eff − v_{e} . hff & = v_{f} . (ehf)_{0}. \end{matrix} \}  (249)

A similar analysis may be applied to effect the elimination of the fourth pair of ratios (238), with results entirely analogous. On the whole then it is found, that the four equations (242) express such connexions between the four pairs of ratios (238), as to satisfy not only the two remaining equations, (245) and (246), of their own groups, (e) and (f), but also the six other equations of the two other groups, (g) and (h), included under type VIII.; namely

\begin{matrix}(g) …  [gefh] = 0, [gfeh] = 0, [ghef] = 0; \\ (h) …  [hefg] = 0, [hfeg] = 0, [hgef] = 0; \end{matrix} \}  (250)

for the first line is satisfied by the ratios (249), and the second by the analogous ratios, which are found in a similar way. Thus all the twelve equations of type VIII. are satisfied, if we satisfy only four suitably selected equations of that type; for example, the equations (242): which was what we proposed to demonstrate.

[36.] The eighty equations of association, assigned in the Third Section, between the twenty-four constants l_{1} … u_{3}, or (efg), (eff), have therefore, by the recent analysis, been ultimately reduced to sixteen; namely the four equations which thus remain from the last type VIII.; and the twelve others which were contained in the type III., established in that earlier Section: and which (as was lately remarked) leave still no fewer than eight constants arbitrary in this theory of associative quines. We may indeed vary in many ways, consistently with the same general theory, and by the assistance of the other recent types VI. and VII., the system of the sixteen equations of condition which are to be satisfied, and the choice of the eight constants which are to be regarded as still remaining arbitrary and undetermined: and it may not be useless, nor uninteresting, to make some remarks hereafter, upon the subject of such selections. But in the mean time it appears to be important to observe, that if some of the recent results, especially the formulæ (210), (228), be combined with some of those previously obtained, and more particularly with the equations (112), (121), of Section III., the following very simple expressions are found, for the ten remaining constants of multplication, the discussion of which had been reserved:

(f) = v_{f}^{2}; (fg) = v_{f} . v_{g};  (251)

or, with the notations abc, and with the usual cyclical permutation of the indices 1, 2, 3,

a_{1} = v_{1}^{2}, a_{4} = v_{4}^{2}, b_{1} = v_{2}v_{3}, c_{1} = v_{1}v_{4}.  (252)

If then we write for abridgement,

\begin{matrix}v & = v_{1}x_{1} + v_{2}x_{2} + v_{3}x_{3} + v_{4}x_{4}, \\ v′ & = v_{1}x′_{1} + v_{2}x′_{2} + v_{3}x′_{3} + v_{4}x′_{4}, \end{matrix} \}  (253)

the square of any quadrinomial vector ϖ, and the scalar of the product of two such vectors, will take these remarkably simple forms:

ϖ^{2} = v^{2}; Sϖϖ′ = v . v′;  (254)

this latter scalar thus decomposing itself into a product* of two linear functions of the constituents, namely those here denoted by v and v′. And because it is easy to prove, from what has been already shown, compare (244), that in the present theory the constants v_{e} are connected by relations of the form

−v_{e} . efe = v_{e} . fee = v_{f} . eff + v_{g} . efg + v_{h} . efh,  (255)

we find, by multiplying this equation by v_{g}, and attending to (251), the following theorems for those general associative quines which have been in this section considered:

\begin{matrix}0 = Sι_{g} Vι_{e}ι_{f} = Sι_{g}ι_{e}ι_{f}; \\ 0 = Sϖϖ′ϖ′′; 0 = (Vϖϖ′)^{2}; \end{matrix} \}  (256)

results which may be compared with some of those obtained in Section IV., for the two particular quine-systems, (A) and (B).

Observatory of Trinity College, Dublin, February 14, 1855.

[To be continued.]

* A similar decomposition into linear factors takes place for the quadrinomes (A) of par. [13.], but at the expense of one of the six arbitrary constants l_{1} l_{2} l_{3} m_{1} m_{2} m_{3}, when we establish between those symbols the relation,

l_{1}m_{1}^{2} + l_{2}m_{2}^{2} + l_{3}m_{3}^{2} = l_{1}l_{2}l_{3} + 2m_{1}m_{2}m_{3}.

In general, I find that it is possible to satisfy all the conditions of association for polynomes, and at the same time to secure a decomposition of Sϖϖ′ into linear factors, while yet preserving so many as 3n − 4 constants of multiplication arbitrary. (For quadrinomes, 3n − 4 = 9 − 4 = 5; for quines 3n − 4 = 12 − 4 = 8.)